### v3

# biblio.bib

Abstract
Directed acyclic graphs (DAGs) are directed graphs in which there is no path from a vertex to itself. DAGs are an omnipresent data structure in computer science and the problem of counting the DAGs of given number of vertices and to sample them uniformly at random has been solved respectively in the 70’s and the 00’s. In this paper, we propose to explore a new variation of this model where DAGs are endowed with an independent ordering of the out-edges of each vertex, thus allowing to model a wide range of existing data structures.
We provide efficient algorithms for sampling objects of this new class, both with or without control on the number of edges, and obtain an asymptotic equivalent of their number. We also show the applicability of our method by providing an effective algorithm for the random generation of classical labelled DAGs with a prescribed number of vertices and edges, based on a similar approach. This is the first known algorithm for sampling labelled DAGs with full control on the number of edges, and it meets a need in terms of applications, that had already been acknowledged in the literature.
1 Introduction
Directed Acyclic Graphs (DAGs for short) are directed graphs in which there is no directed path (sequence of incident edges) from a vertex to itself. They are an omnipresent data structure in various areas of computer science and mathematics. In concurrency theory for instance, scheduling problems usually define a partial order on a number of tasks, which is naturally encoded as DAG via its Hasse diagram [CMPTVW2010, CEH2019]: each task corresponds to a vertex in the graph and task dependencies are materialised by directed edges. Scheduling then corresponds to finding a good topological order on this graph. Natural question such as counting or sampling the schedulings of a program are studied in this context for the purpose of random testing [GS2005]. DAGs also appear as the result of the compression of some tree-like structures such as XML documents [BLMN2015]. In functional programming in particular, this happens at the memory layout level of persistent tree-like values, where the term “hash-consing” has been coined to refer to this compression [Goto1974]. Computer algebra systems also make use of this idea to store their symbolic expression [Ershov1958]. These tree use cases leverage the fact that actual memory gains can be obtained by compacting trees, which has been quantified in [FSS1990] and has motivated the study of compacted trees in the recent years [EFW2021, GW2024]. Finally, complex histories, such as those used in version control systems (see Git for instance [git2018, p. 17]) or genealogy “trees” are DAGs as well.
Most of the applications presented here actually require to add some more structure on the space of DAGs in order to faithfully model the objects at play, which is the main motivation of the present article. We first give some background on the combinatorics of DAGs and then expand on our contributions.
1.1 Background on DAGs
Two different models of DAGs have received a particular interest: labelled DAGs and unlabelled DAGs. The most obvious one is the labelled model, in which one has a set $V$ of vertices (often $\llbracket 1;n\rrbracket$) connected by a set of edges $E\subseteq V\times V$. The term labelled is used because the vertices can be distinguished here, they can be assigned labels. On the other hand, unlabelled DAGs are the quotient set obtained by considering labelled DAGs up to relabelling, that is to say up to a permutation of their vertices (which is reflected on the edges). These two types of objects serve a different purpose, the former represents relations over a given set whereas the latter represents purely structural objects. From a combinatorial point of view, a crucial difference between the two models is that one has to deal with symmetries when enumerating unlabelled DAGs which makes the counting and sampling problem significantly more involved.
Counting
The problem of counting DAGs has been solved in early 70’s by Robinson and Stanley using different approaches. In [Robinson1970], Robinson exhibits a recursive decompositions of labelled DAGs leading to a recurrence satisfied by the numbers $A_{n,k}$ of DAGs with $n$ vertices including $k$ sources (vertices without any incoming edge). He later reformulates those results in terms of a new kind of generating functions, now called graphical generating functions in [Robinson1973], and also obtains the asymptotic number of size $n$ DAGs. Around the same time, Stanley also used a generating function approach in [Stanley1973] obtained the same results by deriving identities of the chromatic polynomial. Robinson also solves the unlabelled case starting from the same ideas but using Burnside’s lemma and cycle index sums to account for the symmetries of these objects. He provides a first solution in [Robinson1970] and makes it more computationally tractable in [Robinson1977]. In the 90’s, Gessel generalised those results, also using the graphical generating function framework in [Gessel1995, Gessel1996] to take into account more parameters and count DAGs by vertices and edges, but also sinks and sources.
Random sampling
From the point of view of uniform random generation, the recursive decomposition exhibited by Robinson in [Robinson1973] is interesting as it is amenable to the recursive method pioneered by Nijenhuis and Wilf in [NW1978]. This yields a polynomial time algorithm for sampling uniform DAGs with $n$ vertices. The analysis of this algorithm has been done in [KM2015] but it had been acknowledged earlier in [MDB2001] although the article proposes an alternative solution. Both [KM2015] and [MDB2001] also offer a Markov chain approach to the random sampling problem and an interesting discussion on the pros and cons of both approaches is given in [KM2015]. Remote from the field of combinatorics, the random generation of DAGs is also an active topic in the area of applied statistics and Bayesian inference. In this context, DAGs encode a relevant structure in a collection of random variables and the problem of interest is to sample DAGs from a particular distribution related to those random variables. To this end, authors resort both to Monte Carlo Markov Chains approaches [KSM2022, KM2017] and methods similar to what is referred to as the recursive method in combinatorics [TVK2020]. An important point in [KM2017] is better performance can be achieved by using a combination of both approaches, in particular by exploiting the combinatorial properties of DAGs. Notable is that sampling from the uniform distribution is tackled as a particular case in [TVK2020] and solved with the asymptotically optimal $O(n^{2})$ complexity at the expense of a $O(n^{3})$ pre-processing step.
Unfortunately, to our knowledge, no efficient uniform random generator of unlabelled has been found yet. Moreover, unlike in the labelled case, the method derived by Robinson to exhibit the number of unlabelled DAGs cannot be easily leveraged into a random sampler as they make extensive use of Burnside’s lemma.
Another interesting question is that of controlling the number of edges in those random samplers. Indeed, sampling a uniform DAG with a prescribed number of vertices and edges cannot be achieved using the Markov chain approach as it constrains the chain too much, and the formulas of Gessel are not amenable to this either. In [KM2015, § 7], the authors provide an interesting discussion on which kind of restrictions can be made on DAGs with the Markov chain approach. They address in particular the case of bounding the number of edges and highlight that the Monte Carlo Markov Chain approach fails when the desired number of edges is too low, thus advocating for having precise combinatorial enumerations.
1.2 Contributions
In the present paper, we propose to study an alternative model of DAGs, which we call Directed Ordered Acyclic Graphs (DOAGs), and which are enriched with additional structure on the edges. More precisely, a DOAG in an unlabelled DAG where (1) set of outgoing edges of each vertex is totally ordered and (2) the sources are totally ordered as well. This local ordering of the outgoing edges allows to capture more precisely the structure of existing mathematical objects. For instance, the compressed formulas and tree-like structures mentioned earlier (see [Ershov1958, Goto1974]) indeed present with an ordering as soon as the underlying tree representation is ordered. This is the case for most trees used in computer science (e.g. red-black trees, B-trees, etc.) and for all formulas involving non-commutative operators. The model we introduce thus allows for a more faithful modelling of a wide range of objects. We present here several results regarding DOAGs, as well as an extension of our method to classical labelled DAGs.
As a first step of our analysis, we describe a recursive decomposition scheme that allows us to study DOAGs using tools from enumerative combinatorics. This allows us to obtain a recurrence formula for counting them, as well as a polynomial-time uniform random sampler, based on the recursive method from [NW1978], giving full control over their number of vertices and edges. Our decomposition is based on a “vertex-by-vertex” approach, that is we remove one vertex at a time and we are able to describe exactly what amount of information is necessary to reconstruct the graph. This differs from the approach of Robinson to study DAGs, where all the sources of a DAG are removed at once instead. Although this is a minor difference, our approach allows us to easily account for the number of edges of the graph, which is why our random sampler is able to target DOAGs with a specific number of edges. In terms of application, this means that we are able to efficiently sample DOAGs of low density. A second by-product of our approach is that it makes straightforward to bound the out-degree of each vertex, thus allowing to sample DOAGs of low degree as well.
In order to show the applicability of our method, we devise a similar decomposition scheme for counting labelled DAGs with any number of vertices, edges, and sources. This allows us to transfer our results on DOAGs in the context of labelled DAGs. Our new recurrence differs from the formula of Gessel [Gessel1996] in that it does not resort to the inclusion-exclusion principle. Our approach allows us to obtain an efficient uniform random sampler of labelled DAGs with a prescribed number of vertices, edges, and sources. Here again, in addition to giving control over the number of edges of the produced objects, our approach can also be adapted to bound the out-degree of their vertices. To our knowledge, this is the first such sampler.
Finally, in a second part of our study of DOAGs, we focus on their asymptotic behaviour and get a first result in this direction. We consider the number $D_{n}$ of DOAGs with $n$ vertices, one source, and any number of edges, and we manage to exhibit an asymptotic equivalent of an uncommon kind:
$D_{n}\sim c\cdot n^{-1/2}\cdot e^{n-1}\prod_{k=1}^{n-1}k!\quad\text{for some constant~{}$c>0$.}$
In the process of proving this equivalent, we state an upper bound on $D_{n}$ by exhibiting a super-set of the set of DOAGs of size $n$, expressed in terms of simple combinatorial objects: variations. This upper-bound is close enough to $D_{n}$ so that we can leverage it into an efficient uniform rejection sampler of DOAGs with $n$ vertices and any number of edges. Combined with an efficient anticipated rejection procedure, allowing to reject invalid objects as soon as possible, this lead us to an asymptotically optimal uniform sampler of DOAGs of size $n$.
In terms of applications, our random generation algorithms enable to experiment with the properties of the objects they model and with the average complexity of algorithms operating on them. A similar approach is for instance taken in [CP2024] where samplers for a realistic class of Git graphs are developed in order to tackle the average complexity of a new algorithm introduced in [Lecoq2024, CDL2022]. Random testing a also an important application of random sampling, especially as a building block for property-based testing, a now well-established framework pioneered Claessen and Hughes in [CH2000].
This paper extends an earlier article [GPV2021] with new results on the asymptotics of DOAGs, with an optimal uniform random sampler for the case when the number of edges is not prescribed, and covers a larger class of DOAGs and DAGs by drooping a constraint on the number of sinks. For the sake of completeness, the most important results and ideas from [GPV2021] will be recalled in the present paper, but the reader will have to refer the earlier article to get the full proof and algorithmic details.
1.3 Outline of the paper
In Section 2, we start by introducing the class of Directed Ordered Acyclic Graphs and their recursive enumeration and describe a recursive decomposition scheme allowing to count them. In Section 3, we quickly go over earlier results regarding the random generation of DOAGs with a prescribed number of vertices, edges, and sources. The presentation given in this paper slightly generalises over the algorithm given in [GPV2021] but the ideas and proofs remain unchanged. Section 4 shows that our approach applies to labelled graphs as well and opens the way for further research regarding this class. We show that our method, when applied to labelled DAGs, yields a constructive counting formula for them, that is amenable to efficient uniform random generation with full control on the number of edges. Then, in Section 5, we present a bijection between DOAGs and class of integer matrices. This bijection is the key result of this paper as it allows to understand the structure of DOAGs in detail, and to obtain both asymptotic and algorithm results in the following sections. In Section 6, we present a first asymptotic result: we give an asymptotic equivalent of the number of DOAGs of size $n$ with any number of sources and edges. We also state some simple structural properties of those DOAGs. In light of the matrix encoding and these asymptotic results, we design an optimal uniform random sampler of DOAGs with a given number of vertices (but no constraint on the number of edges), that is described in Section 7.
An implementation of all the algorithms presented in this paper is available at https://github.com/Kerl13/randdag.
2 Definition and recursive decomposition
In this section, we recall a model of directed acyclic graphs called “Directed Ordered Acyclic Graphs” (or DOAGs) that we introduced in [GPV2021]. It is similar to the classical model of unlabelled DAGs but where, in addition, we have a total order on the outgoing edges of each vertex. The presentation we opted for here slightly differs from that of [GPV2021] but essentially defines the same objects, the only difference being that we now allow several sinks for the sake of generality.
Definition 1 (Directed Ordered Graph).
A directed ordered graph (or DOG for short) is a triple $(V,E,{(\prec_{v})}_{v\,\in V\cup\{\top\}})$ where:
  • $V$ is a finite set of vertices;
  • $E\subset V\times V$ is a set of edges;
  • for all $v\in V$$\prec_{v}$ is a total order over the set of outgoing edges of $v$;
  • and $\prec_{\top}$ is a total order over the set of sources of the graph, that is the vertices without any incoming edge.
Two such graphs are considered to be equal if there exists a bijection between their respective sets of vertices that preserves both the edges and the order relations $\prec_{v}$ and $\prec_{\top}$.
Definition 2 (Directed Ordered Acyclic Graph).
A directed ordered acyclic graph (or DOAG for short) is a directed ordered graph $(V,E,{(\prec_{v})}_{v\,\in V\cup\{\top\}})$ such that $(V,E)$, seen as a directed graph, is acyclic.
We study this class as a whole, however, some sub-classes are also of special interest, in particular for the purpose of modelling compacted data structures. Tree structures representing real data, such as XML documents for instance [BLMN2015], are rooted trees. When these trees are compacted, the presence of a root translates into a unique source in the resulting DOAG. Similarly, DOAGs with a single sink will arise naturally when compacting trees which bear a single type of leaves. In particular the model of compacted binary trees, which can also be seen as a class of cycle-free binary automata, has been shown have unusual combinatorial properties in [EFW2020, EFW2021] and corresponds to a restriction of our model with only binary nodes (and one sink). For these reasons, we will also discuss how to approach the sub-classes of DOAGs with a single source and/or a single sink in this document.
In order to illustrate the definition, the first line of Figure 1 depicts all the DOAGs with at most $3$ vertices and the second line shows all the DOAGs with exactly $4$ vertices and $3$ edges. There are 17 of them while there are $95$ DOAGs with $4$ vertices in total.
Figure 1: All DOAGs with respectly $1$ vertex, $2$ vertices, $3$ vertices, and simultaneously $4$ vertices and $3$ edges. All edges are implicitly oriented from top to bottom, the blue labels and arrows represent the sources and out-edges orderings (always from left to right).
2.1 Recursive decomposition
We describe a canonical way to recursively decompose a DOAG into smaller structures. The idea is to remove vertices one by one in a deterministic order, starting from the smallest source (with respect to their ordering $\prec_{\top}$). Formally, we define a decomposition step as a bijection between the set of DOAGs with at least two vertices and the set of DOAGs given with some extra information.
Let $D$ be a DOAG with at least $2$ vertices and consider the new graph $D^{\prime}$ obtained from $D$ by removing its smallest source $v$ and its outgoing edges. We need to specify the ordering of the sources of $D^{\prime}$. We consider the ordering where the new sources of $D^{\prime}$ (those that have been uncovered by removing $v$) are considered to be in the same order (with respect to each other) as they appear as children of $v$ and all larger than the other sources. The additional information necessary to reconstruct $D$ from $D^{\prime}$ is the following:
  1. 1.
    the number $s\geq 0$ of sources of $D^{\prime}$ which have been uncovered by removing $v$;
  2. 2.
    the (possibly empty) set $I$ of internal (non-sources) vertices of $D^{\prime}$ such that there was an edge in $D$ from $v$ to them;
  3. 3.
    the function $f:I\to\llbracket 1;s+|I|\rrbracket$ identifying the positions, in the list of outgoing edges of $v$, of the edges pointing to an element of $I$.
More formally, this decomposition describes a bijection between DOAGs with at least $2$ vertices and quadruples of the form $(D^{\prime},s,I,f)$ where:
  • $D^{\prime}$ is a DOAG (obtained by removing $v$ from $D$);
  • $I$ is any subset of the internal vertices of $D^{\prime}$ (children of $v$ in $D$);
  • $s$ is any integer between 00 and the number of sources of $D^{\prime}$;
  • and $f:I\to\llbracket 1;s+|I|\rrbracket$ is an injective function (mapping the vertices of $I$ to their positions in the list of children of $v$ in $D$).
In order to prove that this is indeed an bijection, we consider the inverse transformation below. Start with a quadruple $(D^{\prime},s,I,f)$ as described above. Add a new source $v$ in $D^{\prime}$ with $s+|I|$ outgoing edges such that the $i$-th of these edges is connected to $f^{-1}(i)$ when $i\in f(I)$ and is connected to one of the $s$ largest sources of $D^{\prime}$ otherwise. The $s$ largest sources of $D^{\prime}$ must be connected to the new source exactly once and in the same order as they appear in the list of sources of $D^{\prime}$. The resulting graph is a DOAG and it is easy to check that this mapping and the decomposition are inverses of each other.
Note that the order in which the vertices are removed when iterating this process corresponds to a variant of the BFS algorithm where only sources are eligible to be picked next in the search, and their are picked in the order described above. Figure 2 illustrates this decomposition by applying the first two steps on a large example DOAG.
Figure 2: The two first steps of the recursive decomposition of a DOAG by removing sources one by one in a breadth first search (BFS) fashion. The edges are implicitly oriented from top to bottom and the order of the outgoing edges of each vertex is indicated by the thinner blue arrows (always from left to right here). The integer labels at each stage indicate the ordering of the sources. The big disk, square, and triangle are only here to highlight particular vertices involved with the functions $f$ in the decomposition.
This decomposition can be used to establish a recursive formula for counting DOAGs, which is given below. Let $D_{n,m,k}$ denote the number of DOAGs with $n$ vertices, $m$ edges and $k$ sources, then we have:
$D_{1,m,k}$ $=\boldsymbol{1}_{\{m=0\,\land\,k=1\}}$ (1)
$D_{n,m,k}$ =$=\begin{cases}0&\text{when}~{}k\leq 0\\ \sum_{p=0}^{n-k}\sum_{i=0}^{p}D_{n-1,m-p,k-1+p-i}\binom{n-k-p+i}{i}\binom{p}{i}i!&\text{otherwise,}\end{cases}$
where $p=s+i$ corresponds the out-degree of the smallest source. The term $\binom{n-k-p+i}{i}=\binom{n-k-s}{i}$ accounts for the choice of the set $I$ and the term $\binom{p}{i}i{!}$ accounts for the number of injective functions $f:I\to\llbracket 1;p\rrbracket$. The upper bound on $p$ in the sum is justified by the fact that the out-degree of any vertex can be at most the number of non-sources in the graph, that is $(n-k)$.
The decomposition scheme presented here differs from the approach described by Robinson in [Robinson1970] as it operates on only one source at a time. It is also reminiscent of the peeling processes used in map enumeration where maps are decomposed one face at a time, see for instance [KMSW2019]. However, the absence of ordering amongst the incoming edges of each vertex in our setup renders those approaches inapplicable as is.
2.1.1 Special sub-classes based on out-degree constraints
Since $p=i+s$ is the out-degree of the removed source in the above summation, it is easy to adapt this sequence for counting DOAGs with constraints on the out-degree of the vertices. For instance, DOAGs with only one sink are obtained by ensuring that every vertex has out-degree at least one. In other words, let the summation start at $p=1$. Note that restricting DOAGs to have only one single sink or one single source ensures that they remain connected, however not all connected DOAGs are obtained this way. As another example, DOAGs with out-degree bounded by some constant $d$ are obtained by letting $p$ range from 00 to $\min(n-k,d)$.
The general principle is that the DOAGs whose vertices’ out-degrees are constrained to belong to a given set $\mathcal{P}$, are enumerated by the following generalised recurrence.
$D^{\mathcal{P}}_{1,m,k}$ $=\boldsymbol{1}_{\{m=0\,\land\,k=1\}}$ (2)
$D^{\mathcal{P}}_{n,m,k}$ =$=\begin{cases}0&\text{when}~{}k\leq 0\\ \sum_{p\in\mathcal{P}}\sum_{i=0}^{p}D^{\mathcal{P}}_{n-1,m-p,k-1+p-i}\binom{n-k-p+i}{i}\binom{p}{i}i!&\text{otherwise,}\end{cases}$
The first values of the sequence $D_{n,m}=\sum_{k}D_{n,m,k}$ counting DOAGs by number of vertices and edges only are given in Table 1. Table 2 gives the first values of $D^{\mathcal{P}}_{n}=\sum_{m,k}D^{\mathcal{P}}_{n,m,k}$ for some relevant choices of $\mathcal{P}$. None of these sequences seem to appear in the online encyclopedia of integer sequences (OEIS [^1] https://oeis.org/) yet.
Table 1: Number of DOAGs with $n$ vertices and $m$ edges for small values of $n$ and $m$.
$n$ $D_{n}$ $D_{n,m}=\sum_{k}D_{n,m,k}$ for $m=0,1,2,3,\ldots$
$1$ $1$ $1$
$2$ $2$ $1,1$
$3$ $8$ $1,2,3,2$
$4$ $95$ $1,3,8,17,27,27,12$
$5$ $4858$ $1,4,15,48,139,349,718,1136,1272,888,288$
$6$ $1336729$ $1$, $5$, $24$, $100$, $391$, $1434$, $4868$, $14940$, $40261$, $92493$, $175738$, $266898$, $310096$, $258120$, $136800$, $34560$
Table 2: Number of DOAGs with $n$ vertices and a constrained set of allowed degrees.
Restrictions sequence
$\mathcal{P}=\mathbb{N}$
(all DOAGs)
$1$, $2$, $8$, $95$, $4858$, $1336729$, $2307648716$, $28633470321822$, $2891082832793961795$, $2658573971407114263085356$, $24663703371794815015576773905384$, $\ldots$
$\mathcal{P}=\mathbb{N}$, $k=1$
(1 source)
$1$, $1$, $4$, $57$, $3399$, $1026944$, $1875577035$, $24136664716539$, $2499751751065862022$, $2342183655157963146881571$, $22043872387559770578846044961204$, $\ldots$
$\mathcal{P}=\mathbb{N}^{\star}$, $k=1$
(1 source, 1 sink)
$1$, $1$, $3$, $37$, $2103$, $627460$, $1142948173$, $14701782996075$, $1522511169925136833$, $1426529804350999351686869$, $13426022673540053054145359653988$, $\ldots$
$\mathcal{P}=\{0,1,2\}$, $k=1$
(unary-binary)
$1$, $1$, $4$, $23$, $191$, $2106$, $29294$, $495475$, $9915483$, $229898277$, $6074257926$, $180460867600$, $5962588299084$, $\ldots$
3 Earlier results on counting and recursive sampling
In this section we summarise our earlier results on the counting and random sampling problem for DOAGs when all three parameters (number of vertices, edges and sources) are fixed. The theorems are stated in a slightly more general setting here than in [GPV2021] so as to capture all variants of the model as described in Section 2.1.1. However, there is no technical difficulty in the generalisation so that the proofs from [GPV2021] still apply, almost without modification.
We first present a utility result: computing all the numbers $D_{n,m,k}^{\mathcal{P}}$ up to a certain bound on $n$$m$, and $k$ can be done in polynomial time. This is of moderate interest in itself, but this is a requirement for our samplers, that compute theses values as a pre-processing step. Our algorithm is based on the so-called “recursive method” from [NW1978].
3.1 Counting
As mentioned above, tabulating the values of the sequence $D_{n,m,k}^{\mathcal{P}}$ can be done in polynomial time. This means that this counting pre-processing step is tractable up to a certain point.
Theorem 1.
Let $N,M>0$ be two integers. And let $\mathcal{P}$ be a subset of $\mathbb{N}$ such that $\mathcal{P}\cap\llbracket 0;n\rrbracket$ can be enumerated in linear time in $n$. Computing $D_{n,m,k}^{\mathcal{P}}$ for all $n\leq N$, all $m\leq M$, and all possible $k$ can be done with $O(N^{4}M)$ multiplications of integers of size at most $O(\max(M,N)\ln N)$.
The bound given here is independent of $\mathcal{P}$ and thus pessimistic. If $\mathcal{P}$ is bounded (bounded out-degree DOAGs) or sparse, the algorithm will perform better. In practice, the cost of the counting process is actually the limiter factor for the recursive sampler presented below. Indeed, it is hard to reach sizes of the order of the thousands because of the large amount of time and memory necessary to compute and store all the numbers.
3.2 Recursive random sampling
A straightforward application of the recursive method from Nijenhuis and Wilf [NW1978] leads to Algorithm 1, which is presented here in a high level fashion.
Algorithm 1 Recursive uniform sampler of DOAGs
1:Three integers $(n,m,k)$ such that $D_{n,m,k}^{\mathcal{P}}>0$
2:A uniform random DOAG with $n$ vertices (including $k$ sources) and $m$ edges
3:function UnifDOAGP($n,m,k$)
4:     if $n=0$ or $n=1$ then generate the (unique) DOAG with $n$ vertex
5:     else
6:         pick $(p,i)$ with probability ${D_{n-1,m-p,k-1+p-i}^{\mathcal{P}}\binom{n-k-p+i}{i}\binom{p}{i}i{!}}/{D_{n,m,k}^{\mathcal{P}}}$
[…]
8:         $I\leftarrow$ a uniform subset of size $i$ of the inner vertices of $D^{\prime}$
9:         $f\leftarrow\text{a uniform injection from~{}$I$ to~{}$\llbracket 1;p\rrbracket$}$
10:         return $\operatorname{\mathrm{decomp}}^{-1}(D^{\prime},p-i,I,f)$
In [GPV2021, §3], we discuss how to implement Algorithm 1 efficiently. In particular we suggest a data-structure to represent DOAGs that allows for an efficient selection of the subset $I$ at line 8 and the function $f$ at line 9. In addition, implementation considerations are also given for the pick instruction at line 6, which is the core of the “recursive method”. As mentioned above, the numbers $D_{n,m,k}^{\mathcal{P}}$ either have to be pre-computed for Algorithm 1 to work, or must be lazily computed and memoised on the fly.
In practice, pre-computing all the necessary numbers to sample a uniform DOAG with $50$ vertices (without any constraint on $m$ and on the out-degree) using our library already takes about $8$ seconds on a standard laptop. This running time rapidly increases, which makes the cost generating large structures prohibitive. However, when limiting the number of edges and using a finite set $\mathcal{P}$, one can achieve much better results. For instance generating the four large bounded-degree DOAGs from Figure 3(b) takes about $11$ seconds on the same laptop, most of this time being spent in the pre-computation.
Theorem 2.
Algorithm 1 computes a uniform random DOAG with $n$ vertices (among which $k$ are sources) and $m$ edges by performing $O\left(\sum_{v}d_{v}^{2}\right)$ multiplications of a small integer by a large integer, where $v$ ranges over the vertices of the resulting graph and $d_{v}$ is the out-degree of $v$.
Note that the sum $\sum_{v}d_{v}^{2}$ is of the order of $m^{2}$ in the worst case but can be significantly smaller, in particular if $\mathcal{P}$ is bounded or sparse. In the best case we have $d_{v}\sim\frac{m}{n}$ for most of the vertices and as a consequence $\sum_{v}d_{v}^{2}\sim{m^{2}}/{n}$. Also note that in order for the algorithm to be made generic in $\mathcal{P}$, we only have to use the sequence $D_{n,m,k}^{\mathcal{P}}$ rather than $D_{n,m,k}$, which reflects the generality of the recursive method.
Four random DOAGs, drawn using Algorithm 1 with $\mathcal{P}=\{0,1,2\}$$n=1250$$m=1300$ and $k=1$ are pictured in Figure 3(b). As a comparison, a truncated version of the Git history of the linux kernel is pictured on the left in Figure 3(a). Expectedly, the Git history looks more structured. This is because developers work on short-lived branches, consisting of chains of commits, generally starting from the main branch and merged back on main after the new feature (or bug fix, etc.) has been completed, reviewed and accepted.
(a) The Git history of the Linux kernel, truncated at depth $70$, as of the 14th of April 2025, just after the release of version 6.15-rc2. It has $1270$ vertices and $1431$ edges.
(b) Four DOAGs drawn uniformly at random amongst all DOAGs with $1250$ vertices, $1300$ edges and with out-degree bounded by $2$.
Figure 3: The graphical representation of a (truncated) Git history and four bounded-degree random DOAGs.
This explains the chains of unary vertices and the triangular patterns. Although the random DOAGs look more “random”, they exhibit similar smaller-scale structures such as triangular patterns and locally denser areas. It has to be noted that in order to obtain those pictures, we had to reduce the number of edges compared to the Git graph on the left as uniform DOAGs with $n=1270$ and $m=1431$ are already visually too dense to look like a Git graph. We recall that the DOAG model presented in this paper aims at being a general purpose modelling tool and thus does not integrate Git specific constraints. In this regard, a random DOAG is not expected to have all the structural properties of a Git graph. However, the comparison in Figure 3 showcases that the model can be tweaked (here by controlling $m$ and the out-degree) in order to resemble some application-inspired graphs. In the case of Git, we obtain a somewhat similar shape.
4 Extension to labelled DAGs
In this section we demonstrate how our decomposition scheme can be applied to the classical model of labelled DAGs to obtain new recurrences on known sequences. In the 1990s, Gessel [Gessel1996] already obtained equations allowing to count labelled DAGs by vertices, edges, and sources (and also sinks actually) using a generating functions approach. These equations involve the inclusion-exclusion principle which has one drawback: they are usually not amenable to efficient random generation. The reason for this is that subtractions translate into rejections in the recursive algorithm, that are here too costly to be usable. In the present paper, we derive new recurrences with a combinatorial meaning and that do not involve the inclusion-exclusion principle. As a consequence, we can obtain an efficient random sampler of DAGs with full control over the number of vertices, sources, and edges.
As in Section 2, we present here a slight generalisation of the formula given in [GPV2021, §4] allowing to capture various classes of labelled DAGs. We omit the proofs here as they follow a straightforward adaptation of the arguments given in [GPV2021] and the recursive method.
4.1 Recursive decomposition
The key idea to our decomposition is to consider labelled DAGs with a distinguished source (this operation is called pointing) and to decompose them by removing this source. This describes a bijection between source-pointed labelled DAGs and labelled DAGs endowed with some additional structure, in the same fashion as in Section 2 of the present paper.
Let $\mathcal{P}$ be any subset of $\mathbb{N}$ and let $A_{n,m,k}^{\mathcal{P}}$ denote the number of labelled DAGs with $m$ edges and $n$ vertices including $k$ sources, and in which every vertex except the (first) sink has out-degree in $\mathcal{P}$. The number of such DAGs with a distinguished (or pointed) source is given by $k\cdot A_{n,m,k}^{\mathcal{P}}$ since any of the $k$ sources may be distinguished. Let $D$ denote one such DAG and let $v$ denote its distinguished source. Removing the distinguished source in $D$ and decrementing the labels of the vertices of higher label than $v$ by one yields a regular vertex-labelled DAG $D^{\prime}$ with $n-1$ vertices. Moreover, the three pieces of information that are necessary to reconstruct the source are the following:
  1. 1.
    the label $\ell$ of the source $v$ which has been removed;
  2. 2.
    the set $S$ of sources of $D^{\prime}$ which have been uncovered by removing $v$;
  3. 3.
    the set $I$ of internal (non-sources) vertices of $D^{\prime}$ that were pointed at by $v$.
The reconstruction is then straightforward:
  • increment all the labels that are greater or equal to $\ell$ by one;
  • create a new vertex labelled $\ell$ and “mark” it: this is the distinguished source;
  • add edges from $\ell$ to all the vertices from $S$ and $I$.
This decomposition is simpler than that of DOAGs because there is no ordering to maintain here. Hence, any subset $S$ of the set of sources of $D^{\prime}$ is licit here. The triplet $(\ell,S,I)$ is thus not constrained which leads to the simple counting formula, given below, where $p$ denotes the out-degree of $v$ (and thus the cardinality of $S\cup I$).
$A_{1,m,k}^{\mathcal{P}}$ $=\boldsymbol{1}_{\{m=0~{}\wedge~{}k=1\}}$ (3)
$kA_{n,m,k}^{\mathcal{P}}$ =$=\begin{cases}n\sum_{p\in\mathcal{P}\cap\llbracket 0;n-k\rrbracket}\sum_{i=0}^{p}A_{n-1,m-p,k-1+p-i}^{\mathcal{P}}\binom{n-k-p+i}{i}\binom{k-1+p-i}{p-i}&\text{if~{}$1\leq k$}\\ 0&\text{otherwise.}\end{cases}$
In the last equation:
  • the factor $k$ on the left counts the number of ways to choose the distinguished source;
  • the factor $n$ on the right counts the number of ways to choose the label of the new source;
  • and the two binomial coefficient count the number of ways to select the subsets $I$ and $S$.
When $\mathcal{P}=\mathbb{N}$, we recover the sequence counting all labelled DAGs, known as A003024 in the OEIS and first enumerated in [Robinson1970, Stanley1973, Robinson1973]. For $\mathcal{P}=\mathbb{N}^{\star}$ and with $k=1$, we find the number of labelled DAGs with a single source and a single sink, known to Gessel in [Gessel1995, Gessel1996] and stored at A165950 in the OEIS.
4.2 Random generation
A recursive random sampling algorithm similar to Algorithm 1 from Section 3 can be obtained from formula (3). The only difference in methodology from Algorithm 1 is that one has to deal with the marking of the sources here and thus the division by $k$ at the third line of (3). It can be handled as follows: at every recursive call, first generate a labelled DAG with a distinguished source (counted by $k\cdot A_{n,m,k}$) and then forget which source was distinguished. Since the recursive formula for $k\cdot A_{n,m,k}$ has no division, the uniform sampler of marked DAGs is obtained using the standard recursive method. Moreover, forgetting which source was marked does not introduce bias in the distribution since all sources have the same probability to be marked. A uniform random sampler of labelled DAGs with $n$ vertices, $k$ sources, and $m$ edges is described in Algorithm 2.
Algorithm 2 Uniform random sampler of vertex-labelled DAGs.
Three integers $(n,m,k)$ such that $A_{n,m,k}^{\mathcal{P}}>0$
A uniform random labelled DAG with $n$ vertices (including $k$ sources and one sink), $m$ edges, and in which every vertex (except the first sink) has out-degree in $\mathcal{P}$.
function unifDAGP($n,m,k$)
     if $n=0$ or $n=1$ then generate the (unique) labelled DAG with $n$ vertex
     else
         pick $(p,i)$ with probability $\frac{A_{n-1,m-p,k-1+p-i}^{\mathcal{P}}\binom{n-k-p+i}{i}\binom{k-1+p-i}{p-i}}{A_{n,m,k}^{\mathcal{P}}}$
         $D^{\prime}$leftarrow unifDAGP($n-1,m-p,k-1+p-i$)
         $I$leftarrow a uniform subset of size $i$ of the inner vertices of $D^{\prime}$
         $S$leftarrow a uniform subset of size $(p-i)$ the sources of $D^{\prime}$
         $\ell$leftarrow Unif($\llbracket 1;n\rrbracket$)
         relabel $D^{\prime}$ by adding one to all labels $\ell^{\prime}\geq\ell$
         return the DAG obtained by adding a new source to $D^{\prime}$ with label $\ell$ and with an outgoing edge to every vertex of $I\cup S$
5 Matrix encoding
In this section, we introduce the notion of labelled transition matrices and give a bijection between DOAGs and these matrices, thus offering an alternative point of view on DOAGs. These results are key ingredients of the paper, since they enable us, in the next two sections, to prove an asymptotic equivalence for the number of DOAGs with $n$ vertices, and to design a efficient uniform random sampler for those DOAGs. We also recall here the definition and basic properties of variations, which are an elementary combinatorial object playing a central role in our analysis.
5.1 The encoding
The decomposition scheme described in Section 2 corresponds to a traversal of the DOAG. This traversal induces a labelling of the vertices from $1$ to $n$, which allows us to associate the vertices of the graph to these integers in a canonical way. We then consider its transition matrix using these labels as indices. Usually, the transition matrix of a directed graph $D$ is defined as the matrix ${(a_{i,j})}_{1\leq i,j\leq n}$ such that $a_{i,j}$ is $1$ if there is an edge from vertex $i$ to vertex $j$ in $D$, and 00 otherwise. This representation encodes the set of the edges of a DAG but not the edge ordering of DOAGs. In order to take this ordering into account, we use a slightly different encoding.
Definition 3 (Labelled transition matrix of a DOAG).
Let $D$ be a DOAG with $n$ vertices. We associate the vertices of $D$ to the integers from $1$ to $n$ corresponding to their order in the vertex-by-vertex decomposition. The labelled transition matrix of $D$ is the matrix ${(a_{i,j})}_{1\leq i,j\leq n}$ with integer coefficients such that $a_{i,j}=k>0$ if and only if there is an edge from vertex $i$ to vertex $j$ and this edge is the $k$-th outgoing edge of $i$. Otherwise $a_{i,j}=0$.
An example of a DOAG and its transition matrix are pictured in Figure 4. The thick lines are not part of the encoding and their meaning will be explained later when we characterise which integer matrices can be a labelled transition matrix. Let $\phi$ denote the function mapping a DOAG to its labelled transition matrix. This function is clearly injective as the edges of the graph can be recovered as the non-zero entries of the matrix, and the ordering of the outgoing edges of each vertex is given by the values of the corresponding entries in each row. Characterising the image of $\phi$ however requires more work.
Figure 4: An example DOAG and its labelled transition matrix, the zeros are represented by the absence of a number.
We can make some observations. First, by definition of the traversal of the DOAG, the labelled transition matrix of a DOAG is strictly upper triangular. Indeed, since the decomposition algorithm removes one source at a time, the labelling it induces is a topological sorting of the graph. Moreover, since the non-zero entries of row $i$ encode the ordered set of outgoing edges of vertex $i$, these non-zero entries form a permutation. More formally:
  • a non-zero value cannot be repeated within a row;
  • and if a row contains $d\geq 1$ non-zero entries, then these are the integers from $1$ to $d$, in any order.
Informally, these two properties ensure that a matrix encodes a labelled DOAG (a DOAG endowed with a labelling of its vertices) and that this labelling is a topological sorting of the graph. However, they are not enough to ensure that this topological sorting is precisely the one that is induced by the decomposition. The matrices satisfying these two properties will play an important role in the rest of the paper. We call them “variation matrices”.
Definition 4 (Variation).
A variation is a finite sequence of non-negative integers such that
  1. 1.
    each strictly positive number appears at most once;
  2. 2.
    if $0<i<j$ and $j$ appears in the sequence, then $i$ appears too.
The size of a variation is its length.
[…]
Definition 5 (Variation matrix).
Let $n>0$ be a positive integer. A matrix of integers ${(a_{i,j})}_{1\leq i,j\leq n}$ is said to be a variation matrix if
  • it is strictly upper triangular;
  • for all $11\leq i\leq n-$ 1, the sub-row ${(a_{i,j})}_{i<j\leq n}$ is a variation (of size $n-i$).
From an enumerative point of view, a variation matrix can be seen as a sequence of variations $(v_{1},v_{2},\ldots,v_{n-1})$ where for all $11\leq i\leq n-$ 1, the variation $v_{i}$ has size $i$.
We have established that all labelled transition matrices of DOAGs are variation matrices. Note that the converse is not true. For instance, the matrix pictured in Figure 5 is a variation matrix of size $3$ that does not correspond to any DOAG. The property of this matrix which explains why it cannot be the image of a DOAG is pictured in red on the Figure. The rest of this section is devoted to characterising which of those variation matrices can be obtained as the labelled transition matrix of a DOAG.
Figure 5: An example of a matrix of variations that cannot be obtained as a labelled transition matrix of a DOAG. The labelled DOAG that it encodes is not labelled according to the decomposition order.
Consider a DOAG and its labelled transition matrix. Note that in any column $j$, the non-zero entry with the highest index $i$ (that is in the lowest row on the picture with a non-zero element in column $j$) has a special role: it corresponds to the last edge pointing to vertex $j$ when decomposing the DOAG. This is pictured in Figure 6 where we drew, the same DOAG as in Figure 4 and added:
  • on the right (in the matrix): thick red underlines to show the last non-zero entry of each column;
  • on the left (in the graph): thick red decorations on the last incoming edge (in decomposition order) of each vertex.
When a column has no non-zero entry at all, the top line is pictured in thick red instead. This is the case in the three first columns of the matrix in Figure 6. Still in the figure: in order to make those three extra lines correspond to something in the graph, we added an artificial extra source, connected to all other sources (there is a unique way to do this). Those three extra edges are indeed the last incoming edges of the vertices labelled 1, 2, and 3, that naturally correspond to the red part of the three first columns of the matrix. Note that the thick red edges in the graph on the left of Figure 6 form a spanning tree of the graph, and that the labelling induced by the decomposition coincides exactly with the natural BFS order of the tree.
Figure 6: The same example as Figure 4 with extra decorations to highlight the correspondence between the last incoming edge (in decomposition order) of each vertex and the last non-zero entry of the columns of its labelled transition matrix. An artificial $\bot$ vertex, connected to every source, has been added in the graph in order to show that the thick red edges form a spanning tree of the graph.
Another important remark is that, when several underlined cells occur on the same row $i$ in the matrix, they correspond to several sources that are discovered at the same decomposition step of the DOAG (upon removing the same source). Recall that the decomposition
Figure 7: Same-row underlined entries are sorted.
algorithm sorts the labels of these new sources by following the total order of the outgoing edges of vertex $i$. This implies that the underlined entries within the same row have to be increasingly sorted (from left to right). For instance, observe that there are three consecutive underlined cells in the first row of the matrix in Figure 6. Indeed, when removing the first source of the DOAG on the left, we uncover three new sources which are respectively in first, second and fourth position in the outgoing edges order of the removed source.
Figure 8: Thick red lines draw a descending staircase.
A second key property is that if more than one underlined cell occur in a row of the matrix, these are always the first non-zero entries in that row. This is because, the decomposition algorithm consumes sources in the same order in discovers them. As a consequence, for a given vertex $i$, those of its children that become sources upon removing $i$ will be processed before any other children, and thus appear first in the list of the non-zero entries of the $i$-th row. Put differently, the red thick lines drawn in Figure 4 is visually a staircase that only goes down when moving toward the right of the matrix.
The two properties that we just described actually characterise the variation matrices that can be obtained as the labelled transition matrices of a DOAG. This is stated in a more formal manner in Theorem 3.
Theorem 3.
All labelled transition matrices of DOAGs are variation matrices. Furthermore, let $A={(a_{i,j})}_{1\leq i,j\leq n}$ be a variation matrix, and for all $j\in\llbracket 1;n\rrbracket$, let $b_{j}$ denote the largest $i\leq n$ such that $a_{i,j}>0$ if such an index exists and 00 otherwise. Then, $A$ is the labelled transition matrix of some DOAG if and only if the two following properties hold:
  • the sequence $j\mapsto b_{j}$ is weakly increasing;
  • whenever $0<b_{j}=b_{j+1}$, we have that $a_{b_{j},j}<a_{b_{j},j+1}$.
The sequence $(b_{j})_{1\leq j\leq n}$ from the theorem is the formalisation of the thick red lines from Figure 6. The first condition from the theorem corresponds to the ever-descending nature of the “staircase”, as illustrated in Figure 8 The second condition corresponds to the ordering of underlined cells within a row, as illustrated in Figure 7.
Proof.
The fact that the labelled transition matrix of a DOAG is a variation matrix is clear from the definition. We prove the rest of the theorem in two steps.
Step 1: labelled transition matrices satisfy the conditions. Let $D$ be a DOAG of size $n$ and let $A$ be its labelled transition matrix. Let $b={(b_{j})}_{1\leq j\leq n}$ be defined as in the statement of the theorem. We shall prove that it satisfies the two properties of the theorem. The case $n=1$ is trivially and we proceed by induction when $n\geq 2$.
When $n\geq 2$, we can decompose $D$ as a quadruplet $(D^{\prime},s,I,f)$ and we have that the labelled transition matrix of $D^{\prime}$ is the sub-matrix $A^{\prime}$ of $A$ obtained by removing its first row and first column, that is $A^{\prime}=(a_{i+1,j+1})_{1\leq i,j\leq n-1}$. We can define the sequence $b^{\prime}=(b_{j}^{\prime})_{1\leq j\leq n-1}$ corresponding to $A^{\prime}$ similarly to the sequence $b$.
We distinguish between three cases.
  • If $j$ is such that $b_{j}=0$, then $b_{j}\leq b_{j+1}$ automatically, and there is no second condition to check.
  • If $j$ is such that $b_{j}=1$, then $b_{j+1}$ cannot be zero, otherwise that would mean that the vertex labelled $(j+1)$ is a source of $D$ but the vertex labelled $j$ is not. Indeed, since the sources of $D$ are processed before any other vertex by the decomposition algorithm, they get the smallest labels. Hence $b_{j}\leq b_{j+1}$.
    In addition, if $b_{j}=b_{j+1}=1$, then the vertices labelled $j$ and $(j+1)$ both become sources, at the same time, upon removing the first vertex. By construction of the decomposition, they get labels in an order compatible with the order of the outgoing edges of the first source, and thus we have $a_{1,j}<a_{1,j+1}$.
  • Finally, if $j$ is such that $b_{j}\geq 2$, then we have $b_{j}=b_{j-1}^{\prime}+1$. By induction we also have that $b_{j}^{\prime}\leq b_{j-1}^{\prime}$, which, in particular, implies that there is at least one non-zero entry in the $j$-th column of $A^{\prime}$ and thus in the $(j+1)$-th column of $A$. It follows that $b_{j+1}=b_{j}^{\prime}+1$ and finally $b_{j}\leq b_{j+1}$ and $b_{j}=b_{j+1}\implies a_{b_{j},j}<a_{b_{j},j+1}$ by induction.
Step 2: any matrix satisfying the conditions is a labelled transition matrix. Let $A$ be a variation matrix of size $n$ and let $b$ be as in the statement of the theorem and satisfying the two given properties. We shall prove that $A$ is the image by $\phi$ of some DOAG.
Let $V=\llbracket 1;n\rrbracket$ and $E=\left\{(i,j)\in\llbracket 1;n\rrbracket^{2}~{}|~{}a_{i,j}>0\right\}$. We have that $(V,E)$ defines an acyclic graph since $A$ is strictly upper-triangular. In addition, for each $v\in V$, define $\prec_{v}$ to be the total order on the outgoing edges of $v$ in $(V,E)$ such that $u\prec_{v}u^{\prime}$ if and only if $a_{v,u}<a_{v,u^{\prime}}$ in $A$. This is well defined since the outgoing edges of $v$ are precisely the integers $j$ such that $a_{v,j}>0$ and since the non-zero entries of the row $v$ are all different by definition of variation matrices. Finally, define $\prec_{\top}$ to be the total order on the sources of $(V,E)$ such that $u\prec_{\top}v$ if and only if $u<v$ as integers. Let $D$ be the DOAG given by $(V,E,{(\prec_{v})}_{v\,\in V\cup\{\top\}})$.
Remember that DOAGs are considered up to a permutation of their vertices that preserves $E$ and $\prec$. In order to finish this proof, we have to check that the particular labelling encoded by $V$ is indeed the labelling induced by the decomposition of $D$. Then it will be clear that $\phi(D)=A$ and we will thus have exhibited a pre-image of $A$.
First, since $A$ is strictly upper-triangular, its first column contains only zeros and thus $1$ is necessarily a source of $D$. In addition, by definition of $\prec_{\top}$, it must be the smallest source. Then, upon removing $i$, one of two things can happen:
  • either $D$ has more than one source, in which case $2$ is the second source by monotony of the sequence ${(b_{j})}_{1\leq j\leq n}$;
  • or $1$ was the unique source of $D$, in which case the next source to be processed is its first child. The children of $1$ are the integers $j$ such that $b_{j}=1$. By monotony of $b_{j}$ again (or triangularity of the matrix), $2$ is necessarily a child of $1$. Moreover, by the second property of the sequence $b$, we have that for all $j<j^{\prime}$ such that $b_{j}=b_{j^{\prime}}=1$$a_{1,j}<a_{1,j^{\prime}}$.
In both case, we proved that $2$ is the second vertex to be processed. We can then repeat this argument on the DOAG obtained by removing $1$, which corresponds to the matrix ${(a_{i,j})}_{2\leq i,j\leq n}$ and conclude by induction.∎
We have now established that the encoding $\phi$ of DOAGs as labelled transition matrices is a bijection from DOAGs to the matrices described in Theorem 3. From now on, we will write “a labelled transition matrix” to refer to such a matrix. We can also state a few simple properties of these matrices. By definition we have that
  • the number of vertices of a DOAG is the dimension of its labelled transition matrix;
  • the number of edges of a DOAG is the number of non-zero entries of the matrix;
  • the sinks of the DOAG correspond to the zero-filled rows of the matrix;
  • the sources of the DOAG correspond to the zero-filled columns of the matrix.
Furthermore, the first property of the sequence ${(b_{j})}_{1\leq j\leq n}$ defined in Theorem 3 implies that the zero-filled columns of the matrix must be contiguous and on the left of the matrix. The number of sources of the DOAG is thus the maximum $j$ such that column $j$ is filled with zeros.
We will see in the next section that working at the level of the labelled transition matrices, rather than at the level of the graphs, is more handy to exhibit asymptotic behaviours. This will also inspire an efficient uniform random sampler of DOAGs with $n$ vertices in Section 7.
6 Asymptotic results
The characterisation of the labelled transition matrices of DOAGs gives a more global point of view on them compared to the decomposition given earlier, which only looks locally around one source. By approaching the counting problem from the point of matrices, we manage to provide lower and upper bounds on the number of DOAGs with $n$ vertices (and any number of edges). These bounds are precise enough to give a good intuition on the asymptotic behaviour of these objects, and we then manage to refine them into an asymptotic equivalent for their cardinality. Building on this same approach, we provide in Section 7 an efficient uniform sampler of DOAGs with $n$ vertices.
This section is mostly devoted to proving Theorem 4. Sub-sections 6.1 to 6.3 present the general approach and give all the intermediate results that are necessary to prove Theorem 4. We conclude this section by giving asymptotic estimations of two relevant parameters of DOAGs under the uniform model: their number of sources and edges. We obtain these two last results by leveraging the work done in this section on the matrix point of view on DOAGs.
Theorem 4.
There exists a constant $c>0$ such that the number $D_{n}$ of DOAGs with $n$ vertices and the number $D_{n}^{\star}$ of such DOAGs having only one source satisfy
[…]
$\sim c\cdot e^{\zeta^{\prime}(-1)-1}\cdot n^{-7/12}\cdot{\big{(}e\sqrt{2\pi}\big{)}}^{n}\!\cdot e^{-\frac{3}{4}n^{2}}\cdot n^{n^{2}/2}$
[…]
The super factorial provides a concise way to express this equivalent and also reflects the relation between DOAGs and variation matrices, which will be further developed in this section.
6.1 First bounds on the number of DOAGs with $\boldsymbol{n}$ vertices
Let $D_{n}=\sum_{m,k}D_{n,m,k}$ denote the number of DOAGs with $n$ vertices and any number of sources and edges. By Theorem 3, all labelled transition matrices are variation matrices. A straightforward upper bound for $D_{n}$ is thus given by the number of variation matrices of size $n$.
Lemma 1 (Upper bound on the number of DOAGs).
For all $n\geq 1$, the number $D_{n}$ of DOAGs of size $n$ satisfies
[…]
Lemma 2 (Asymptotic estimation of the super-factorial [Barnes1900]).
When $n\to\infty$, we have
[…]
where $\zeta$ denotes the Riemann zeta function.
In order to prove Lemma 1, we first need to give estimates for the number $v_{n}$ of variations of size $n$.
Lemma 3 (Exact and asymptotic number of variations).
For all $0\leq p\leq n$, the number $v_{n}$ of variations of size $n$, and the number $v_{n,p}$ of variations of size $n$ containing exactly $p$ zeros, are respectively given by
$v_{n}=n!\sum_{j=0}^{n}\frac{1}{j!}\quad\text{and}\quad v_{n,p}=\frac{n!}{p!}$
As a consequence $v_{n}\leq e\cdot n{!}$ and $v_{n}=e\cdot n!+o(1)$.
Proof.
Let $0\leq p\leq n$, a variation of size $n$ containing exactly $p$ zeros is the interleaving of a permutation of size $(n-p)$ with an array of zeros of size $p$. As a consequence
$v_{n,p}=\binom{n}{p}(n-p)!=\frac{n!}{p!}.$
We then get $v_{n}$ and the asymptotic estimate by summation:
$v_{n}=\sum_{p=0}^{n}v_{n,p}=n!\sum_{p=0}^{n}\frac{1}{p!}=n!\left(\sum_{p=0}^{\infty}\frac{1}{p!}-\sum_{p=n+1}^{\infty}\frac{1}{p!}\right)=en!-\sum_{p=n+1}^{\infty}\frac{n!}{p!},$
which allows to conclude since the last sum is $\sum_{p>n}\frac{n!}{p!}=O(n^{-1})$. ∎
The proof of Lemma 1 follows from this lemma.
Proof of Lemma 1.
By inclusion, there are less DOAGs of size $n$ that there are variation matrices. In addition, a variation matrix is given by a sequence $v_{1},v_{2},\ldots,v_{n-1}$ of variations such that for all $i$$v_{i}$ is of size $i$. We thus have the following upper bound for $D_{n}$:
[…]
Figure 9: Lower-bound on the set of labelled transition matrices.
Obtaining a lower bound on $D_{n}$ requires to find a subset of the possible labelled transition matrices described in Theorem 3 that is both easy to count and large enough to capture a large proportion of the DOAGs. One possible such set is that of the labelled transition matrices which have no zero values on the super-diagonal ${(a_{i,i+1})}_{1\leq i<n}$. These matrices are picture in Figure 9 on the right.
These correspond to DOAGs such that, at every step of the decomposition, we have only one source and thus uncover exactly one new source. In such matrices, the properties of the sequence ${(b_{j})}_{1\leq j\leq n}$ from Theorem 3 are automatically satisfied. Intuitively, forcing the super-diagonal to be positive still leaves a large amount of free space on the right of that diagonal to encode many possible DOAGs, so it should be expected that it gives a decent lower bound.
Lemma 4 (A first lower bound on the number of DOAGs).
There exists a constant $A>0$ such that for all $n\geq 1$, we have
[…]
Proof.
In a labelled transition matrix with positive values on the super-diagonal, the $i$-th row can be seen as a variation of size $(n-i)$ that does not start with a zero. Moreover, the number of variations of size $n$ starting with a zero is actually the number of variations of size $(n-1)$ so that the number of possibilities for the $i$-th row of the matrix we count here is $(v_{n-i}-v_{n-i-1})$. In addition, by Lemma 3, we also have that
$v_{n}-v_{n-1}=e\cdot n!-e\cdot(n-1)!+o(1)=e\cdot n!\cdot\frac{n-1}{n}\left(1+O\left(\frac{1}{n!}\right)\right).$ (4)
Note that when $i=n-1$, we have $v_{n-i}-v_{n-i-1}=v_{1}-v_{0}=1$. Indeed, the row of index $i=n-1$ contains only the number $1$ in the super diagonal since at the last step of the decomposition, we have two connected vertices and there is only one such DOAG. Setting aside this special case, which does not contribute to the product, we get the following lower bound for $D_{n}$:
[…]
where the first product telescopes and yields $\frac{1}{n-1}$ and the second one converges to a constant as $n\to\infty$. This allows to conclude the proof the lemma. ∎
[…]
Lemma 5 (A better lower bound for the number of DOAGs).
There exists a constant $B>0$ such that, for all $n\geq 1$, we have
[…]
Proof.
In order to obtain this lower bound, we count the number of valid labelled transition matrices such that all but exactly one of the cells on the super-diagonal have non-zero values. Furthermore, in order to avoid having to deal with border cases, we assume that the unique zero value on the super-diagonal appears between $i=2$ and $i=n-2$. Let $22\leq i\leq n-$ 2 and let us consider those matrices ${(a_{p,q})}_{1\leq p,q\leq n}$ such that $a_{i,i+1}=0$. Those matrices are illustrated in Figure 10
Figure 10: Illustration of the matrices with only one zero on the super-diagonal
The differences between these matrices and those enumerated in the proof of the previous lemma are the following (assuming $i$ is that unique index such that $a_{i,i+1}=0$).
  1. 1.
    On row $i-1$, the two first cells on the right of the diagonal ($a_{i-1,i}$ and $a_{i-1,i+1}$) must have positive values and must be in increasing order. In the case of $a_{i-1,i}$, this is because it is on the super diagonal. And for $a_{i-1,i+1}$, this is because $a_{i,i+1}=0$: since it is above this cell, and since the cell $a_{i-1,1}$ on its left is non-zero, it must be non-zero. Otherwise the condition from Theorem 3 are violated. In terms of DOAGs, this means that the vertex $i-1$ produces two new sources when it is removed but vertex $i$ produces none.
  2. 2.
    On row $i$, any variations of size $(n-i)$ starting by a zero is allowed.
We get the number of variations of size $n$ starting by two increasing positive values (condition 1 above) by inclusion-exclusion. That is,
  • consider all the variations of size $n$ ($v_{n}$ possibilities);
  • remove the number of variations that have a zero in first position ($v_{n-1}$ possibilities);
  • remove the number of variations that have a zero in second position ($v_{n-1}$ possibilities);
  • add the number of variations that start with two zeros, because they have been removed twice in the two previous lines ($v_{n-2}$ possibilities);
  • and finally, divide by two because only half of these matrices have their first values in increasing order.
This yields the following formula for counting such variations:
$\frac{v_{n}-2v_{n-1}+v_{n-2}}{2}\underset{n\to\infty}{\sim}\frac{e}{2}\cdot n!.$
As a consequence, the total number of labelled transition matrices considered at the beginning of the proof, such that $a_{i,i+1}=0$, is given by
$\frac{v_{n-i-1}-2v_{n-2-i}+v_{n-3-i}}{2}\cdot v_{n-i-1}\cdot\prod_{\begin{subarray}{c}1\leq p\leq n-1\\ p\not\in\{i-1,i\}\end{subarray}}(v_{n-p}-v_{n-p-1})$
$=\frac{(v_{n-i-1}-2v_{n-2-i}+v_{n-3-i})v_{n-i-1}}{2(v_{n-i+1}-v_{n-i})(v_{n-i}-v_{n-i-1})}\cdot\prod_{p=1}^{n-1}(v_{n-p}-v_{n-p-1}).$
By summing over $22\leq i\leq n-$ 2, we get
$\sum_{i=2}^{n-2}\frac{(v_{n-i-1}-2v_{n-2-i}+v_{n-3-i})v_{n-i-1}}{2(v_{n-i+1}-v_{n-i})(v_{n-i}-v_{n-i-1})}\cdot\prod_{p=1}^{n-1}(v_{n-p}-v_{n-p-1}).$ (6)
The fraction in the last equation is equivalent to $\frac{1}{2(n-i)}$ when $n-i\to\infty$. So after a change of variable, the above sum is equivalent to
$\sum_{i=2}^{n-2}\frac{(v_{i-1}-2v_{i-2}+v_{i-3})v_{i-1}}{2(v_{i+1}-v_{i})(v_{i}-v_{i-1})}\sim\sum_{i=2}^{n-2}\frac{1}{2i}\sim\frac{\ln(n)}{2}$
[…]
6.2 Obtaining the polynomial term by bootstrapping
Let us denote $P_{n}$ the polynomial term in $D_{n}$, that is the quantity
[…]
We have proved above that for some constant $B>0$, we have $B\frac{\ln(n)}{n}\leq P_{n}\leq 1$. A consequence of these inequalities is that for all $k\in\mathbb{Z}$, we have
$P_{n+k}\leq 1\leq\frac{n}{B\ln(n)}P_{n}\underset{n\to\infty}{=}o(nP_{n}).$ (7)
Note that we did the extra work in Lemma 5 in order to get the extra $\ln(n)$ factor that is crucial to get the $o$ term. Equation (7) allows to justify that $P_{n+k}/n$ is negligible compared to $P_{n}$, for any constant values of $k$. Although intuitively the contrary would be surprising, this fact is not clear a priori as an arbitrary polynomial sequence $P_{n}$ could have violent oscillations for some values of $n$. This is a key ingredient for proving an asymptotic equivalent for $P_{n}$.
To refine our knowledge on the sequence $P_{n}$, we use a decomposition of the labelled transition matrices focused on the values it takes near the diagonal on its first rows. We categorise the possible labelled transition matrices ${(a_{i,j})}_{1\leq i,j\leq n}$ into the four following cases.
Case 1: $\boldsymbol{a_{1,2}=0}$.
In this case, the first source is not connected to the second vertex and the matrix has thus more than one source. The first row of such a matrix is a variation of size $(n-2)$ and the lower part ${(a_{i,j})}_{2\leq i,j\leq n}$ encodes a DOAG of size $(n-1)$, the DOAG obtained by removing the first source. However, it is important to note that not all combinations of a size-$(n-2)$ variation and a size-$(n-1)$ matrix yield a valid size-$n$ labelled transition matrix. For instance, a variation of the form $v=(0,1,0,2,\ldots)$ and a lower matrix with at least three sources cannot be obtained together as they would violate the constraints of Theorem 3.
Case 2: $\boldsymbol{a_{1,2}>0\wedge a_{2,3}>0}$.
In this case, the first row is a variation of size $(n-1)$ starting by a positive value, and the lower part ${(a_{i,j})}_{2\leq i,j\leq n}$ encodes a DOAG of size $(n-1)$ with exactly one source, again obtained by removing the first source. This second fact is a direct consequence of $a_{2,3}>0$. Here, all such pairs can be obtained.
Case 3: $\boldsymbol{a_{1,2}>0\wedge a_{2,3}=0\wedge a_{3,4}>0}$.
In this case the lower part ${(a_{i,j})}_{3\leq i,j\leq n}$ encodes a DOAG of size $n-2$ with exactly one source, the first row is necessarily a variation of size $(n-1)$, starting with two positive increasing values, and the second row is a variation of size $(n-2)$ starting by a zero. Here again this decomposition is exact: all such triplets can be obtained here.
Case 4: $\boldsymbol{a_{1,2}>0\wedge a_{2,3}=a_{3,4}=0}$.
Finally, this case captures all the remaining matrices. The first row is a variation of size $(n-1)$, the second and third rows are variations of sizes $(n-2)$ and $(n-3)$ starting with a zero, and the lower part ${(a_{i,j})}_{4\leq i,j\leq n}$ encodes a size-$(n-3)$ DOAG. Of course, not all such quadruples can be obtained, but this over-approximation will be enough for our proof.
This decomposition into four different cases is illustrated in Figure 11 where $\mathcal{D}$ represents the set of all possible DOAG labelled transition matrices and ${\mathcal{D}}^{\star}$ represents all of those matrices that encode a single-source DOAG.
Figure 11: Decomposition of DOAG labelled transition matrices based or their content near the top of the diagonal. The symbols $\mathcal{D}$ and ${\mathcal{D}}^{\star}$ respectively represent the set of all possible DOAG labelled transition matrices the set of all of those matrices such that $a_{1,2}>0$. The stars ()($\star$)( ⋆ ) represent strictly positive values.
We compute the contributions to $D_{n}$ coming from each of these four terms described above. Let us denote by $D_{n}^{\star}$ the number of DOAG of size $n$ with exactly one source, or equivalently the number of DOAG labelled transition matrices containing a non-zero value at coordinates $(1,2)$. The first line of Figure 11 illustrates the first point of the decomposition, which yields
$D_{n}=D_{n}^{\star}+O(v_{n-2}D_{n-1}).$ (8)
Note that the big-$O$ term comes from the fact that not all pairs made of a size-$(n-2)$ variation and a size-$(n-1)$ labelled transition matrix can be obtained this way, as discussed in the first case above. We could actually have written $0\leq D_{n}-D_{n}^{\star}\leq v_{n-2}D_{n-1}$.
Then we decompose the matrices from ${\mathcal{D}}^{\star}$ depending of their values on the diagonal (cases 2 to 4). The second line of Figure 11 illustrates this decomposition. This translates into the following identity
$D_{n}^{\star}=(v_{n-1}-v_{n-2})D_{n-1}^{\star}+\frac{v_{n-1}-2v_{n-2}+v_{n-3}}{2}v_{n-3}D_{n-2}^{\star}+O\!\left(v_{n-1}v_{n-3}v_{n-4}D_{n-3}\right).$ (9)
[…]
$P_{n}^{\star}=P_{n}+O\!\left(\frac{v_{n-2}}{e(n-1)!}P_{n-1}\right)=P_{n}+O\!\left(\frac{P_{n-1}}{n}\right)=P_{n}+o\!\left(P_{n}\right).$
[…]
$P_{n}^{\star}=\left(1-\frac{1}{n}+O\!\left(\frac{1}{n^{2}}\right)\right)P_{n-1}^{\star}+\frac{1}{2n}\left(1+O\!\left(\frac{1}{n}\right)\right)P_{n-2}^{\star}+O\!\left(\frac{P_{n-3}}{n^{2}}\right)\cdot$ (10)
Since $P_{n}^{\star}\sim P_{n}$ and by equation (7), we have that $O\!\left(P_{n-3}n^{-2}\right)=o\!\left(P_{n}^{\star}n^{-1}\right)$ and that the first term of equation (10) dominates all the others. As a consequence we get a refinement on our knowledge on $P_{n}^{\star}$ (and thus $P_{n}$), that is:
$P_{n}^{\star}\sim P_{n-1}^{\star}.$
It is worth noting that this is the key property that makes analysing $P_{n}^{\star}$ possible. From now on, we know that $P_{n}^{\star}$ does not oscillate, and this is all because of equation (7). By re-using this new information in equation (10), we get another term of the expansion of $P_{n}^{\star}$:
$P_{n}^{\star}=P_{n-1}^{\star}\left(1-\frac{1}{2n}+O\!\left(\frac{1}{n^{2}}\right)\right).$
We conclude on the asymptotic behaviour of $P_{n}^{\star}$ using the following classical argument. The series of general term $\ln\left(\!\frac{P_{n}^{\star}}{P_{n-1}^{\star}}\!\right)+\frac{1}{2n}=O\!\left(n^{-2}\right)$ (defined for $n\geq 2$) is convergent and, if $\lambda$ denotes its sum, we have that
$\lambda-\sum_{j=2}^{n}\left(\ln\left(\!\frac{P_{j}^{\star}}{P_{j-1}^{\star}}\!\right)+\frac{1}{2j}\right)=O(n^{-1}).$
Furthermore, since $P_{1}^{\star}=1$, we also have that
$\sum_{j=2}^{n}\left(\ln\left(\!\frac{P_{j}^{\star}}{P_{j-1}^{\star}}\!\right)+\frac{1}{2j}\right)=\ln P_{n}^{\star}+\frac{1}{2}\left(\ln(n)+\gamma+O\!\left(n^{-1}\right)\right)$
where $\gamma$ denotes the Euler–Mascheroni constant. As a consequence, we have that
$\ln P_{n}^{\star}+\frac{\ln(n)}{2}=\lambda+\gamma+O\!\left(n^{-1}\right)$
and thus
$P_{n}^{\star}=\frac{e^{\lambda+\gamma}}{\sqrt{n}}\left(1+O\!\left(n^{-1}\right)\right).$
By equation (8), we also get that $P_{n}=P_{n}^{\star}\left(1+O\!\left(n^{-1}\right)\right)$, which concludes the proof of Theorem 4, state on page 4 at the beginning of this section.
6.3 Approximation of the constant
Before concluding this section with an analysis of the behaviour of the relevant parameters of DOAGs under the uniform model in sub-Section 6.4, we take a brief detour here to show how to estimate numerically the value of the constant $c$ from Theorem 4.
Let $D_{n,k}$ denote the number of DOAGs with $n$ vertices (including $k$ sources and one sink) and any number of edges. Using the same decomposition as in Section 2 and applying the same combinatorial arguments we get
$D_{n,k}=\sum_{i+s\leq n-k}D_{n-1,k-1+s}\binom{s+i}{s}\binom{n-k-s}{i}i!=\sum_{s\geq 0}D_{n-1,k-1+s}\cdot\gamma(n-k-s,s)$ (11)
where
$\gamma(a,b)=\sum_{i=0}^{a}\binom{b+i}{b}\binom{a}{i}i!.$ (12)
The above sum gives an explicit way to compute $\gamma$, but there is a computationally more efficient way to do so using recursion and memoisation:
$\gamma(a,b)$ $=0$ when $a<0$ or $b<0$ (13)
$\gamma(0,b)$ $=1$ when $b\geq 0$
$\gamma(a,b)$ $=\gamma(a,b-1)+a\cdot\gamma(a-1,b)+\boldsymbol{1}_{\{b=0\}}$ otherwise.
Using this recurrence formula with memoisation, the numbers $D_{n,k}$ for all $n,k\leq N$ can be computed in $O(N^{3})$ arithmetic operations on big integers, which is more efficient than using the recurrence from (1) directly. This is expected because we eliminated the parameter $m$.
[…]
05010015020025000.20.40.60.81
2002102202302402500.490.4950.50.5050.51
[…]
6.4 Asymptotic behaviour of some parameters
We conclude our quantitative study of DOAGs with asymptotic estimations of their typical numbers of sources and edges under the uniform model. The method we apply to get these results builds naturally from the methodology developed in the rest of this section, hence illustrating the usefulness of the matrix-based approach.
6.4.1 Number of sources
It follows from the previous section that the probability that a uniform DOAG of size $n$ has more than one source tends to zero as $n\to\infty$. We can refine this result and compute the probability of having a constant number $k$ of sources.
We have that the number of sources of a DOAG is also the number of empty columns in its labelled transition matrix, and that these columns are necessarily in first positions. Moreover, Theorem 4 gives us the intuition that most of those transition matrices contain positive numbers near the top-left corner of the matrix. We thus split the set of matrices ${(a_{i,j})}_{1\leq i,j\leq n}$ encoding DOAGs with $k$ sources in two categories.
Case $\boldsymbol{a_{k,k+1}>0}$.
Intuitively, the most common scenario is that there is a positive entry in $a_{k,k+1}$. In this case the sub-matrix ${(a_{i,j})}_{k\leq i,j\leq n}$ can be re-interpreted as a DOAG with only one source. Indeed, the condition $a_{k,k+1}$ means that upon removing the $(k-1)$ first sources of the DOAG, the decomposition algorithm does not produce any new source, leaving us with a single-source DOAG. We can characterise those matrices: they are made of $(k-1)$ variations of size $(n-k)$ in the first rows, and a size-$(n-k+1)$ labelled transition matrix corresponding to a single source DOAG below them. Any combination of $(k-1)$ such variations and such a matrix can be obtained.
Case $\boldsymbol{a_{k,k+1}=0}$.
On the other hand we have the matrices such that $a_{k,k+1}=0$. In this case, the $(k-1)$ first rows of the matrix are still variations of size $(n-k)$. The lower part ${(a_{i,j})}_{k\leq i,j\leq n}$ can be seen as a DOAG with at least two sources because its first two columns are empty. Note that here, depending of the $(k-1)$ top variations, we may have restriction on which DOAGs may appear in the lower part. For instance, if $k=2$ and the first row is $(0,0,1,0,0,\ldots,0)$, then the DOAG of size $(n-1)$ without any edge cannot appear in the lower part.
This dichotomy is pictured in Figure 13.
Figure 13: Decomposition of the matrices corresponding to DOAGs with $k$ sources
From the above case analysis, we have the following bounds:
$v_{n-k}^{k-1}D_{n-k+1}^{\star}\leq D_{n,k}\leq v_{n-k}^{k-1}D_{n-k+1}+O\!\left(v_{n-k}^{k-1}(D_{n-k+1}-D_{n-k+1}^{\star})\right).$
Thus, by virtue of Theorem 4, we have the following estimates when $(n-k)\to\infty$
$D_{n,k}=v_{n-k}^{k-1}D_{n-k+1}^{\star}\left(1+O\!\left(\frac{1}{n-k}\right)\right),$ (14)
which allow us to state the following result.
Theorem 5 (Number of sources of uniform DOAGs).
When $n\to\infty$ and $(n-k)\to\infty$, we have that
$D_{n,k}-v_{n-k}^{k-1}D_{n-k+1}=o(D_{n,k})$
where the little oh is uniform: it is arbitrarily smaller than $D_{n,k}$ when $(n-k)\to\infty$. In particular for $k$ constant, we have
$\frac{D_{n,k}}{D_{n}}\sim n^{-\binom{k}{2}}.$
Proof.
The first statement has already been established in equation (14) and the second one is straightforward to obtain using the equivalent $v_{n}\sim en{!}$ for the number of variations. ∎
6.4.2 Number of edges
Another quantity of interests of uniform DOAGs (and graphs in general) is their number of edges. Whereas uniform labelled DAGs have $\frac{n^{2}}{4}$ edges in average, we show here that the number of edges of uniform DOAGs is close to $\frac{n^{2}}{2}$. This has to be compared with their maximum possible number of edges which is $\binom{n}{2}=\frac{n(n-1)}{2}$. This makes uniform DOAGs quite dense objects. The intuition behind this fact is that variations have typically few zeros in them. Indeed, the expected number of zeros of a uniform variation is given by
$\frac{1}{v_{n}}\sum_{p=0}^{n}pv_{n,p}=\frac{n!}{v_{n}}\sum_{p=0}^{n}\frac{p}{p!}=\frac{n!}{v_{n}}\sum_{p=1}^{n}\frac{1}{(p-1)!}\underset{n\to\infty}{\to}1$
where $v_{n,p}$ is the number of variations of size $n$ having exactly $p$ zeros, which is equal to $n!/p{!}$ by Lemma 3. Moreover, the tail of their probability distribution is more than exponentially small:
$\mathbb{P}_{n}[\text{nb zeros}\geq q]=\frac{n!}{v_{n}}\sum_{p=q}^{n}\frac{1}{p!}\underset{n,q\to\infty}{=}\frac{e^{-1}}{q!}\left(1+o\!\left(\frac{1}{n!}\right)\right)\left(1+O\!\left(\frac{1}{q}\right)\right),$
where the first error term depends only on $n$ and the second depends on $q$ and is uniform in $n$. In fact, it is straightforward to show that the distribution of the number of zeros in a uniform variation of size $n$ converges in distribution to a Poisson law of parameter $1$.
Now, recall that DOAG labelled transition matrices are a sub-class of variation matrices, and that the number of non-zero entries in these matrices corresponds to the number of edges of the graph. The above discussion should make the following result intuitive.
Theorem 6 (Number of edges of uniform DOAGs).
The number of edges of a uniform DOAG of size $n$ is, in expectation,
$\binom{n}{2}-O(n).$
Proof.
In terms of labelled transition matrices, the theorem translates into: there is at most a linear number of zeros strictly above the diagonal in the matrix. This is what we prove here.
For all integer $p\geq 0$, by inclusion, we have that the number of DOAG labelled transition matrices with exactly $p$ zeros strictly above the diagonal is upper-bounded by the number $\operatorname{\mathrm{VM}}_{n,p}$ of variation matrices with the same property. Moreover, given a size-$(n-1)$ vector $(p_{1},p_{2},\ldots,p_{n-1})$ of non-negative integers such that for all $i$$p_{i}\leq i$, the number of such variation matrices with exactly $p_{n-i}$ zeros in the $i$-th line is
[…]
By summation over all such vectors such that $\sum_{i=1}^{n-1}p_{i}=p$, we get an expression for $\operatorname{\mathrm{VM}}_{n,p}$:
[…]
In the first sum we have the constraint $p_{i}\leq i$ because a variation has at most $i$ zeros. The inequality comes from the fact that we added more terms in the sum by dropping this constraints. This allows us to interpret the sum as a Cauchy product and can express it as the $p$-th coefficient of the power series $e^{x}\cdot e^{x}\cdot\cdots\cdot e^{x}=e^{(n-1)x}$. It follows that
[…]
As a consequence, we have the following bound for the probability that a uniform DOAG of size $n$ has at most $\binom{n}{2}-q$ zeros:
[…]
The sum in the last equation is the remainder in the Taylor expansion of order $(q-1)$ of the function $e^{x}$ near zero, evaluated at the point $(n-1)$. By using the integral form of this remainder, we have that
$\sum_{p\geq q}\frac{{(n-1)}^{p}}{p!}=\int_{0}^{n-1}e^{t}\frac{{(n-1-t)}^{q-1}}{(q-1)!}dt\leq e^{n-1}\int_{0}^{n-1}\frac{{(n-1-t)}^{q-1}}{(q-1)!}dt=e^{n-1}\frac{{(n-1)}^{q}}{q!}\cdot$
Furthermore, by setting $q=\lambda(n-1)$ for some constant $\lambda>0$, and by using Stirling’s formula, we get that
$\frac{{(n-1)}^{q}}{q!}\sim{\left(\frac{e(n-1)}{q}\right)}^{q}\frac{1}{\sqrt{2\pi q}}\sim{\left(\frac{e^{\lambda}}{\lambda^{\lambda}}\right)}^{n-1}\frac{1}{\sqrt{2\pi\lambda n}}\cdot$
Finally, by using this estimate inside equation (15), and by using Theorem 4 for estimating $D_{n}$, we get that there exists a constant $c^{\prime}>0$ such that
[…]
$\leq\frac{c^{\prime}}{\sqrt{\lambda}}{\left(\frac{e^{\lambda}}{\lambda^{\lambda}}\right)}^{\mathclap{\,\,n-1}}\cdot$
The latter expression is exponentially small as soon as $\lambda>e$ and dominates the tail of the probability distribution of the number of zeros strictly above the diagonal in DOAG labelled transition matrices, which allows to conclude. ∎
7 Uniform sampling of DOAGs by vertices only
The knowledge from the previous section on the asymptotic number of DOAGs with $n$ vertices can be interpreted combinatorially to devise an efficient uniform random sampler of DOAGs based on rejection. Since the set of labelled transition matrices of size $n$ is included in the set of variation matrices of size $n$, a possible approach to sample uniform DOAGs is to sample uniform variation matrices until they satisfy the properties of Theorem 3, and thus encode a DOAG.
Since the number of variation matrices is close (up to a factor of the order of $\sqrt{n}$) to the number of DOAGs, the probability that a uniform variation matrix of size $n$ corresponds to the labelled transition matrix of DOAG is of the order of $n^{-\frac{1}{2}}$. As a consequence, the expected number of rejections done by the procedure outlined above is of the order of $\sqrt{n}$ and its overall cost is $\sqrt{n}$ times the cost of generating one variation matrix. Moreover, we will see that variations (and thus variation matrices) are cheap to sample, which makes this procedure efficient.
This idea, which is a textbook application of the rejection principle, already yields a reasonably efficient sampler of DOAGs. In particular it is much faster than the sampler from the previous section based on the recursive method, because it does not have to carry arithmetic operations on big integers. In this section we show that this idea can be pushed further using “early rejection”. That is to say we check the conditions from Theorem 3 on the fly when generating the variation matrix, in order to be able to abort the generation as soon as possible if the matrix is to be rejected. We will describe how to generate as few elements of the matrix as possible to decide whether to reject it or not, so as to mitigate the cost of these rejections.
First, we design an asymptotically optimal uniform sampler of variations in Section 7.1, and then we show in Section 7.2 how to leverage this into an asymptotically optimal sampler of DOAGs.
7.1 Generating variations
The first key step towards generating DOAGs, is to describe an efficient uniform random sampler of variations. We observe that the number of zeros of a uniform variation of size $n$ obeys a Poisson law of parameter $1$ conditioned to be at most $n$. Indeed,
$\mathbb{P}[\text{a uniform variation of size~{}$n$ has~{}$p$ zeros}]=\frac{v_{n,p}}{v_{n}}\propto\frac{\boldsymbol{1}_{\{0\leq p\leq n\}}}{p!}$
by Lemma 3. A possible way to generate a uniform variation is thus to draw a Poisson variable $p$ of parameter $1$ conditioned to be at most $n$, and then shuffling a size $p$ array of zeros concatenated with the identity permutation using the Fisher-Yates algorithm [FY1948]. This is described in Algorithm 3.
Algorithm 3 Uniform random sampler of variations based on the rejection principle.
1:An integer $n>0$
2:A uniform random variation of size $n$
3:function UnifVariation($n$)
4:     $p$leftarrow BoundedPoisson($1,n$)
5:     $A$leftarrow $[0,0,\ldots,0,1,2,\ldots,n-p]$$\triangleright$ array of length $n$, starting with $p$ zeros
6:     for $i=0$ to $n-2$ do
7:         $r$leftarrow Unif($\llbracket i;n-1\rrbracket$)
8:         $A[r]\leftrightarrow A[i]$triangleright Swap entries of indices $r$ and $i$
9:     return $A$
Regarding the generation of the bounded Poisson variable (performed at line 4), an efficient approach is to generate regular (unbounded) Poisson variables until a value less than $n$ is found. Indeed, the probability $p_{n}$ that a Poisson variable of parameter $1$ is smaller than $n$ is
$p_{n}=e^{-1}\sum_{k=0}^{n}\frac{1}{k!}\geq\frac{2}{e}\qquad\text{when~{}$n\geq 1$}.$
Moreover, when $n$ is large we have $1-p_{n}\sim\tfrac{1}{(n+1)!}$. As a consequence, the expected number of tries of a rejection procedure for sampling conditioned $\text{Poisson}(1)$ variables is $\frac{1}{p_{n}}\leq\frac{e}{2}$. The algorithm described by Knuth in [Knuth1997, page 137] is suitable for our use-case since our Poisson parameter ($1$ here) is small. Furthermore it can be adapted to stop early when values strictly larger than $n$ are found. This is described in Algorithm 4.
Algorithm 4 Adapted Knuth’s algorithm for bounded Poisson simulation
1:A Poisson parameter $\lambda>0$ and an integer $n\geq 0$
2:A Poisson variable of parameter $\lambda$ conditioned to be at most $n$
3:function BoundedPoisson($\lambda,n$)
4:     repeat
5:         $k$leftarrow 00
6:         $p$leftarrow Unif($[0;1]$)
7:         while $(k\leq n)\wedge(p>e^{-\lambda})$ do
8:              $k$leftarrow $k+1$
[…]
10:     until $k\leq n$
11:     return $k$
NB. The $\textsc{Unif}([0;1])$ function generates a uniform real number in the $[0;1]$ interval.
Note that this algorithm relies on real numbers arithmetic. In practice, approximating these numbers by IEEE 754 floating points numbers [ieee754] should introduce an acceptably small error. Indeed, since we only compute products (no sums or subtractions), which generally have few terms, the probability that they introduce an error should not be too far from $2^{-53}$ on a 64-bits architecture. Of course this is only a heuristic argument. A rigorous implementation must keep track of these errors. One possible way would be to use fixed points arithmetic for storing $p$ and to lazily generate the base 2 expansions of the uniform variables at play until we have enough bits to decide how $p$ and $e^{-\lambda}$ compare at line 7. Another way would be to use Ball arithmetic [Hoeven2010, arb] and to increase precision every time the comparison requires more bits. The proofs of correctness and complexity below obviously assume such an implementation.
Lemma 6 (Correctness of Algorithm 3).
Given an input $n>0$, Algorithm 3 produces a uniform random variation of size $n$.
Proof.
The correctness of Algorithm 4 follows from the arguments given in [Knuth1997, page 137], which we do not recall here. Regarding Algorithm 3, the for loop at line 6 implements the Fisher-Yates [FY1948] algorithm, which performs a uniform permutation of the contents of the array independently of its contents. In our use-case, this implies that:
  • the number of zeros is left unchanged;
  • given an initial array with $p$ zeros as shown at line 5, the probability to get a particular variation with $p$ zeros is given by the probability that a uniform permutations maps its first $p$ values to a prescribed subset of size $p$, that is $\frac{p!}{n!}$.
This tells us that, the probability that Algorithm 3 yields a particular variation with $p$ zeros is
$\mathbb{P}[\textsc{BoundedPoisson}(1,n)=p]\cdot\frac{p!}{n!}=\frac{1}{p!\sum_{k=0}^{n}\frac{1}{k!}}\cdot\frac{p!}{n!}=\frac{1}{v_{n}}.\qed$ ∎
The minimal “amount of randomness” that is necessary to simulate a probability distribution is given by its entropy. This gives us a lower bound on the complexity (in terms of random bit consumption) of random generation algorithms. For uniform random generation, this takes a simple form since the entropy of a uniform variable that can take $M$ distinct values is $\log_{2}(M)$. This tells us that we need at least $\log_{2}(v_{n})$ random bits to generate a uniform variation of size $n$. When $n$ is large, we have $\log_{2}(v_{n})=n\log_{2}(n)-\frac{n}{\ln(2)}+O(\log_{2}(n))$. The uniform variation sampler we give in Algorithm 3 is asymptotically optimal in terms of random bit consumption: in expectation, the number of random bits that it uses is equivalent to $\log_{2}(v_{n})$.
Lemma 7 (Complexity of Algorithm 3).
In expectation, Algorithm 3 performs a linear number of arithmetic operations and memory accesses, and consumes $n\log_{2}(n)+o(n\log_{2}(n))$ random bits.
Proof.
The mean of a Poisson variable of parameter $1$ being $1$, Algorithm 4 succeeds to find a value smaller or equal to $n$ in a constant number of tries in average, and each try requires a constant number of uniform variables in average. Furthermore, in order to perform the comparison $p>e^{-1}$ at line 7 in the algorithm, we need to evaluate these uniform random variables. This can be done lazily, and again, it is sufficient to know a constant number of bits of these variables in average to decide whether $p>e^{-1}$.
Regarding the shuffling happening at line 6 in Algorithm 3, it needs to draw $(n-1)$ uniform integers, respectively smaller or equal to 111122223333, …, $n-1$. At the first order, this incurs a total cost in terms of random bits, of
$\sum_{k=2}^{n}\log_{2}(k)\sim n\log_{2}(n).$
In total, the cost of Algorithm 3 is thus dominated by the shuffling, which allows to conclude on its random bits complexity.
Regarding the number of arithmetic operations and memory accesses, generating Poisson variables performs in constant time using similar arguments. The shuffling part of the algorithm is clearly linear. ∎
Note that we count integer operations in the above Lemma, thus abstracting away the cost of these operations. At the bit level an extra $\log_{2}(n)$ term would appear to take into account the size of these integers. This type of considerations is especially important when working with big integers as it was the case in Section 2. However here, arithmetic operations on integers, rather than bits, seems to be the right level of granularity as a real-life implementation is unlikely to overflow a machine integer.
7.2 A fast rejection procedure
Equipped with the variation sampler described above, we can now generate variation matrices in an asymptotically optimal way, by filling them with variations of sizes $(n-1),(n-2),\ldots,3,2,1$. By checking afterwards whether the matrix corresponds to a valid DOAGs, and trying again if not, we get a uniform sampler of DOAGs that is only sub-optimal by a factor of the order of $\sqrt{n}$. This is presented in Algorithm 5. This algorithm is already more efficient than a sampler based on the recursive method, whilst naive.
Algorithm 5 A simple but sub-optimal uniform random sampler of DOAGs
An integer $n>0$
A uniform DOAG with $n$ vertices as its labelled transition matrix
function UnifDOAGNaive($n$)
     $A={(a_{i,j})}_{1\leq i,j\leq n}$leftarrow a zero-filled $n\times n$ matrix
     repeat
         for $i$ from $1$ to $n-1$ do
              ${(a_{i,j})}_{i<j\leq n}$leftarrow UnifVariation($n-i$)
     until $A$ encodes a DOAG
     return The DOAG corresponding to $A$
Checking the validity of a matrix at line 8 corresponds to checking the conditions given in Theorem 3 at page 3. We do not provide an algorithm for this here, as the goal of this section is to iterate upon Algorithm 5 to provide a faster algorithm and get rid of the $\sqrt{n}$ factor in its cost. We will see in the following that checking these conditions can be done in linear time.
Remark. Note that the memory footprint of a DOAG is of the order of $n^{2}$ since it typically has $\frac{n^{2}}{2}$ edges as shown in the previous section, so the number of edges might be a more natural notion of size for those objects when talking about complexity. If we express the complexity of Algorithm 5 in terms of the number of edges $m$ of the generating DOAG, we get that it performs $O(m\sqrt[4]{m})$ memory accesses and consumes $O(m\sqrt[4]{m}\log n)$ random bits. Under this lens, the extra $\sqrt{n}$ factor incurred by the rejection is actually only a fourth root of the more natural size parameter $m$.
As we can see in Theorem 3, the conditions that a variation matrix must satisfy to be a labelled transition matrix concern the shape of the boundary between the zero-filled region between the diagonal and the first positive values above the diagonal. Moreover, we have seen in Theorem 6 that uniform DOAGs tend to have close to $\binom{n}{2}$ edges and thus only a linear number of zeros above the diagonal of their labelled transition matrix. We can thus expect that the area that we have to examine to have access to this boundary should be small. This heuristic argument, hints at a more sparing algorithm that would start by filling the matrix near the diagonal and check its validity early, before generating the content of the whole matrix. This idea, of performing rejection as soon as possible in the generation process, is usually referred to as “anticipated rejection” and also appears in [DFLS2004] and [BPS1994] for instance.
To put this idea in practice, we need to implement lazy variation generation, to be able to make progress in the generation of each line independently, and to perform the checks of Theorem 3 with as little information as necessary.
Ingredient one: lazy variations
Fortunately, Algorithm 3 can be easily adapted for this purpose thanks to the fact that the for loop that implements the shuffle progresses from left to right in the array. So a first ingredient of our optimised sampler is the following setup for lazy generation:
  • for each row of the matrix (i.e. each variation to be sampled), we draw a Poisson variable $p_{i}$ of parameter $1$ and bounded by $(n-i)$;
  • drawing the number at position $(i,j)$, once we have drawn all the numbers of lower coordinate in the same row, can be done by selecting uniformly at random a cell with higher or equal coordinate on the same row and swapping their contents.
This is illustrated in Figure 14.
Figure 14: Ingredient one of the fast rejection-based algorithm: variations can be lazily generated. In the example, the three first elements of the variation at row $i$ are known. When we need to generate its fourth element, we perform a swap of $a_{i,i+4}$ with a uniform cell of index $j\geq i+4$.
Ingredient two: only one initialisation
A straightforward adaptation of Algorithm 3 unfortunately requires to re-initialise the rows after having drawn the Poisson variable (see line 5 of Algorithm 3) at each iteration of the rejection algorithm. This is costly since about $n^{2}/2$ numbers have to be reset. It is actually possible to avoid this by initialising all the rows only once and without any zeros. Only at the end of the algorithm, once a full matrix have been generated, one can re-interpret the $p_{i}$ largest numbers of row $i$, for all $i$, to be zeros. This is pictured in Figure 15.
Figure 15: Ingredient two of the fast rejection-based algorithm: the zeros of the matrix need not be explicitly written. Instead, we interpret the numbers strictly larger than $(n-i-p_{i})$ as zeros. In this example $(n-i)=10$ and $p_{i}=3$ so the numbers 88889999, and $10$ are seen as zeros.
Ingredient three: column by column checking
The last detail that we need to explain is how to check the conditions of Theorem 3. As a reminder
  • for each $2\leq j\leq n$ we need to compute the number $b_{j}=\max\,\{i\ |\ a_{i,j}>0\}$ (or 00 if this set is empty);
  • we must check whether this sequence is weakly increasing;
  • and whenever $b_{j+1}=b_{j}$, we must check that $a_{b_{j},j}<a_{b_{j},j+1}$.
A way of implementing this is to start filling each column of the matrix from bottom to top, starting from the column $j=1$ and ending at column $j=n$. For each column, we stop as soon as either a non-zero number is found or the constraints from Theorem 3 are violated. In order to check these constraints, while filling column $j$ from bottom $(i=j-1)$ to top, we halt as soon as either the cell on the left of the current cell, or the current cell is non-zero. The case when the left cell is non-zero corresponds to when $i=b_{j-1}$ and the conditions of Theorem 3 can be checked. Recall that, per the previous point, the zero test in row $i$ is actually $x\mapsto x>n-i-p_{i}$. We shall prove that this process uncovers only a linear number of cells of the matrix, thus allowing to reject invalid matrices in linear expected time. This idea is pictured in Figure 16.
Figure 16: Ingredient three of the fast rejection-based algorithm: the exploration process of the cells of the matrix is dictated by the following algorithm. We proceed column by column, from bottom to top, and we advance to the next columns as soon as we discover a non-zero cell, or see one on our left. In the pictures, the bullet $\bullet$ represent the current cell, the grey area represents the cells that have not yet been drawn and the thick red lines underline the lowest non-zero cell of each column, as before. Depending on the value that is drawn in the current cell, we either move up or to the next column. Whenever a non-zero cell is on our left, we can decide whether to reject or to keep generating.
The algorithm
Putting all of this together yields Algorithm 6 (on page 6) to generate a uniform DOAG labelled transition matrix using anticipated rejection. The algorithm is split into two parts. First, the repeat-until loop between lines 7 and 22 implements the anticipated rejection phase. At each iteration of this loop, we “forget” what has been done in the previous iterations, so that $A$ is an arbitrary matrix satisfying the following two conditions
$i\geq j\implies a_{i,j}=0$ (16)
$\forall~{}1\leq i<n,\quad\{a_{i,j}\,|\,i<j\leq n\}=\llbracket 1;n-i\rrbracket.$ (17)
The contents of the ${(p)}_{1\leq i<n}$ vector is also forgotten and each value is to be drawn again before any access. The while loop at line 10 implements the traversal of the matrix described above: at each step, the value of the $a_{i,j}$ is drawn and the conditions of Theorem 3 are checked before proceeding to the next step. The array ${(s_{i})}_{1\leq i\leq n}$ stores the state of each lazy variation generator: $s_{i}$ contains the value of the largest $j$ such that $a_{i,j}$ has been drawn. The second part of the algorithm, starting from line 23, completes the generation of the matrix once its near-diagonal part is known and we know no rejection is possible any more. This includes replacing some values of the matrix by 00 because of ingredient two above.
Algorithm 6 An optimised uniform random sampler of DOAGs based on anticipated rejection
1:An integer $n>0$
2:A uniform DOAG with $n$ vertices, encoded as its labelled transition matrix.
3:function UnifDOAGFast($n$)
4:     $A={(a_{i,j})}_{1\leq i,j\leq n}$leftarrow the strictly upper triangular matrix ${(\boldsymbol{1}_{\{j>i\}}\cdot(j-i))}_{1\leq i,j\leq n}$
5:     ${(p_{i})}_{1\leq i<n}$leftarrow uninitialised array
6:     ${(s_{i})}_{1\leq i<n}$leftarrow uninitialised array
7:     $\triangleright$ Anticipated rejection phase
8:         $(i,j)$leftarrow $(1,2)$$\triangleright$ position of the current cell
9:         $p_{1}$leftarrow BoundedPoisson($n-1$)
10:         while $j\leq n$ do
11:              $r$leftarrow Unif($\llbracket j;n\rrbracket$)
12:              $a_{i,r}\leftrightarrow a_{i,j}$
13:              $s_{i}$leftarrow $j$
14:              if $(a_{i,j-1}\leq n-i-p_{i})\wedge(a_{i,j}\not\in\llbracket a_{i,j-1}+1;n-i-p_{i}\rrbracket)$ then
15:                  $\triangleright$ Rejection
16:              else if $a_{i,j}\leq n-i-p_{i}$ then
17:                  $j$leftarrow $j+1$
18:                  $i$leftarrow $j-1$
19:                  $p_{i}$leftarrow BoundedPoisson($1,n-i$)
20:              else
21:                  $i$leftarrow $i-1$
22:     until $j>n$
23:     for $i=1$ to $n-2$ $\triangleright$ Completion of the matrix
24:         for $j=i+1$ to $s_{i}$ do
25:              if $a_{i,j}>n-i-p_{i}$ then $a_{i,j}\leftarrow 0$
26:         for $j=s_{i}+1$ to $n$ do
27:              $r$leftarrow Unif($\llbracket j;n\rrbracket$)
28:              $a_{i,r}\leftrightarrow a_{i,j}$
29:              if $a_{i,j}>n-i-p_{i}$ then $a_{i,j}\leftarrow 0$
30:     return $A$
Lemma 8 (Correction of Algorithm 6).
Algorithm 6 terminates with probability $1$ and returns a uniform random DOAG labelled transition matrix.
This result is a consequence of Algorithm 5 and Algorithm 6 implementing the exact same operations, only in a different order and with an earlier rejection in the latter algorithm. The key characteristic of this new algorithm is that is only needs to perform a linear number of swaps in average to decide whether the reject the matrix or not. As a consequence it is asymptotically optimal in terms of random bits consumption and it only performs about $n^{2}/2$ swaps to fill the $n\times n$ upper triangular matrix.
Theorem 7 (Complexity of Algorithm 6).
In average, in order to generate a uniform DOAG with $n$ vertices, Algorithm 6 performs $\frac{n^{2}}{2}+O(n^{3/2})$ swaps in the matrix, and consumes $\frac{n^{2}}{2}\log_{2}(n)+O(n^{3/2}\log_{2}(n))$ random bits.
Proof.
In the rejection phase, in each column, we draw a certain number of zeros and at most one non-zero value before deciding whether to reject the matrix or to proceed to the next column. As a consequence, when lazily generating a variation matrix, we see at most $(n-1)$ non-zero values and a certain number of zeros that we can trivially upper-bound by the total number of zeros (strictly above the diagonal) in the matrix.
The number of variations of size $n$ with exactly $p$ zeros (with $0\leq p\leq n$) is given by $\frac{n!}{p!}$ by Lemma 3. As a consequence, the expected number of zeros of a variation is given by
$\sum_{p=0}^{n}p\cdot\frac{n!}{p!}\cdot\frac{1}{v_{n}}=\left(e^{-1}+o\!\left(\frac{1}{n!}\right)\right)\sum_{p=0}^{n-1}\frac{1}{p!}=1+O\!\left(\frac{1}{n!}\right).$
It follows that the expectation of the total number of zeros of variation matrix of size $n$ is $n+O(1)$. This proves the key fact that, in expectation, we only discover a linear number of cells of the matrix in the repeat-until loop. Since, in expectation, we only perform $O(\sqrt{n})$ iteration of this loop, it follows that we only perform $O(n^{3/2})$ swaps there. Moreover, one swap costs $O(\log_{2}(n))$ random bits, which thus accounts for a total of $n^{3/2}\log_{2}(n)$ random bits in this loop.
In order to complete the proof, it remains to show that the for loops at the end of Algorithm 6 contribute to the leading terms of the estimates given in the Theorem. The first inner for loop at line 24 replaces, among the already discovered values, the zeros encoded by numbers above the $n-i-p_{i}$ threshold by actual zeros. It is worth mentioning that this only accounts for linear number of operations in total, spanned over several iteration of the outer loop (at line 23). The second inner for loop at line 26 completes the generation of the matrix. The total number of swaps that it performs (and thus the number of uniform variables it draws) is $\frac{n(n-1)}{2}$ minus the number of already discovered cells, that is $n^{2}/2+O(n)$. This allows to conclude the proof. ∎
By Theorem 4 on page 4 in the previous section, we have that $\log_{2}(D_{n})\sim\frac{n^{2}}{2}\log_{2}(n)$. This shows that Algorithm 6 is asymptotically optimal in terms of random bit consumption. Moreover, filling a $n\times n$ matrix requires a quadratic number of memory writes and the actual number of memory access made by our algorithm is of this order too.
8 Conclusion and perspectives
In this paper, we have studied the new class of directed ordered acyclic graphs, which are directed acyclic graphs endowed with an ordering of the out-edges of each of their vertices. We have provided a recursive decomposition of DOAGs that is amenable to the effective random sampling of DOAGs with a prescribed number of vertices, edges and source using the recursive method from Nijenhuis and Wilf. Using a bijection with a class of integer matrices, we also have provided an equivalent for the number of DOAGs with $n$ vertices and designed a uniform random sampler for DOAGs with $n$ vertices and any number of edges. This second sampler is asymptotically optimal, both in terms of memory accesses and random bits consumption.
We have also showed that our approach allows to approach classical labelled DAGs and have obtained a new recurrence formula for their enumeration. The important particularity of this new formula is that it is amenable to effective random sampling when the number of edges is prescribed, which was not the case for previously known formulas.
Perspectives
On DOAGs
So far, we have only approached DOAGs via enumerative tools and ad. hoc. asymptotic techniques. A common and powerful tool in combinatorics is the use of generating function to tackle not only asymptotic estimation, but also the convergence in distribution of parameters, such as the number of edges. The book [FS2009] is a reference on this topic. Classical approaches using ordinary, exponential, or even graphical ([PD2019]) generating functions fail in our context due to the super factorial behaviour of the number of DOAGs. It remains an open question whether it is possible to design a generating function approach to our objects, which could help obtaining finer estimates, not only over $D_{n}$, but also over the low of the number of edges.
Multi-graph variant
An interesting question that is left open by our work is the case of the multi-graph variant of this model: what happens if multiple edges are allowed between two given vertices? This makes the analysis more challenging since there is now an infinite number of objects with $n$ vertices. We thus must change our point of view and take the number $m$ of edges into account in addition to, or instead of, the number of vertices. Estimating the number and behaviour of DOAGs as well as their multi-graph counterpart, when both parameters $n$ and $m$ grow remains an open question and will certainly yield very different results depending on how $n$ and $m$ grow in relation to each other. We argue that the model of multi-edge DOAGs is natural, maybe even more so than that of DOAGs, since they encode (partially) compacted plane trees. Quantitative aspects of tree compaction, in particular the typical compression rate, has been studied in the past [FSS1990] in a general setting. However, the dual point of view that consists in studying already compacted structures directly is a more recent topic, see [EFW2021] and [GW2024] for instance. The class of multi-edge DOAGs generalises the classes studied in those two papers. Moreover, being able to sample them efficiently would give a tool to reach every possible case (including those with double edges) when testing programs manipulating compacted trees (such as compilers) via random generation.
Another interesting question is that of the connectivity. We do not provide a way to count connected DOAGs directly here. However we have already proved that, in the uniform model, they are connected with high probability since they have only one source with high probability. Moreover, since $D_{n}$ grows extremely fast, we can also foresee that a uniform DOAG of size $n$ with two connected components will typically have one tiny component of size $1$ and a big component of size $n-1$, and that the asymptotic estimations of such graphs is straightforward. This implies that sampling a uniform connected DOAG with $n$ vertices is already possible, and efficient, by rejection and the question of their direct enumeration is thus mostly of mathematical interest.
Classical labelled DAGs
Finally, it is also natural to wonder whether our successful approach at efficiently sampling DOAG applies to labelled DAGs. Of course, the asymptotics of DAGs is known [Robinson1973]. But if a matrix encoding similar to ours is feasible, that is an encoding whose combinatorial properties are understood well enough to avoid introduction any bias, then it might be possible to devise an efficient, pre-computation-free, uniform sampler for DAGs.
A starting point in this direction is the fact that uniform labelled DAGs have, in average, $\frac{n^{2}}{4}$ edges. In terms of (upper triangular) adjacency matrices, this means that about half the cells in the upper part of the matrix are non-zero. This is, in a sense, much less dense than for DOAGs. However, this is still dense enough in the sense that the DAG analogue of the red thick path in our figures (described by the sequence $(b_{j})_{1\leq j\leq n}$ in Sections 5 and 7) can be expected to stay close to the diagonal too. As a consequence, the approach proposed in Section 7 for the random generation of DOAGs is still applicable, provided we have an efficient way to sample those paths. Indeed, our fast-rejection procedure in Algorithm 6 can be seen as the combination of two algorithms:
  1. 1.
    an algorithm to sample the path $(b_{j})_{1\leq j\leq n}$ under the distribution induced by DOAGs;
  2. 2.
    and the filling of the remaining cells of the matrix by completing the random variation in reach row.
In order to design a similar approach for DAGs, we need a way to sample the $(b_{j})_{1\leq j\leq n}$ paths (induced by the uniform distribution on DAGs) and the second step of the algorithm would be to fill the rest of the matrix with Bernoulli random variables of parameter $\frac{1}{2}$. Our recent ongoing work on this topic suggests that those paths can indeed be sampled efficiently, which will be investigated further in the near future.
Notable is that the approach presented in [KM2015] can also be seen as a way to work with matrices while maintaining uniformity by using a combinatorial encoding using ordered integer partitions. A caveat however is that their approach still requires a costly pre-processing, which we seek to avoid using a rejection-based approach.
[…]

### v4

@@
-biblio.bib
+Asymptotic analysis and efficient random sampling of directed ordered acyclic graphs
 Abstract
@@
 They are an omnipresent data structure in various areas of computer science and mathematics.
-In concurrency theory for instance, scheduling problems usually define a partial order on a number of tasks, which is naturally encoded as DAG via its Hasse diagram [CMPTVW2010, CEH2019]: each task corresponds to a vertex in the graph and task dependencies are materialised by directed edges.
+In concurrency theory for instance, scheduling problems usually define a partial order on a number of tasks, which is naturally encoded as DAG via its Hasse diagram [6, 4]: each task corresponds to a vertex in the graph and task dependencies are materialised by directed edges.
 Scheduling then corresponds to finding a good topological order on this graph.
-Natural question such as counting or sampling the schedulings of a program are studied in this context for the purpose of random testing [GS2005].
-DAGs also appear as the result of the compression of some tree-like structures such as XML documents [BLMN2015].
-In functional programming in particular, this happens at the memory layout level of persistent tree-like values, where the term “hash-consing” has been coined to refer to this compression [Goto1974].
-Computer algebra systems also make use of this idea to store their symbolic expression [Ershov1958].
-These tree use cases leverage the fact that actual memory gains can be obtained by compacting trees, which has been quantified in [FSS1990] and has motivated the study of compacted trees in the recent years [EFW2021, GW2024].
-Finally, complex histories, such as those used in version control systems (see Git for instance [git2018, p.
+Natural question such as counting or sampling the schedulings of a program are studied in this context for the purpose of random testing [23].
+DAGs also appear as the result of the compression of some tree-like structures such as XML documents [3].
+In functional programming in particular, this happens at the memory layout level of persistent tree-like values, where the term “hash-consing” has been coined to refer to this compression [22].
+Computer algebra systems also make use of this idea to store their symbolic expression [13].
+These tree use cases leverage the fact that actual memory gains can be obtained by compacting trees, which has been quantified in [16] and has motivated the study of compacted trees in the recent years [12, 21].
+Finally, complex histories, such as those used in version control systems (see Git for instance [17, p.
 17]) or genealogy “trees” are DAGs as well.
@@
 The problem of counting DAGs has been solved in early 70’s by Robinson and Stanley using different approaches.
-In [Robinson1970], Robinson exhibits a recursive decompositions of labelled DAGs leading to a recurrence satisfied by the numbers $A_{n,k}$ of DAGs with $n$ vertices including $k$ sources (vertices without any incoming edge).
-He later reformulates those results in terms of a new kind of generating functions, now called graphical generating functions in [Robinson1973], and also obtains the asymptotic number of size $n$ DAGs.
-Around the same time, Stanley also used a generating function approach in [Stanley1973] obtained the same results by deriving identities of the chromatic polynomial.
+In [38], Robinson exhibits a recursive decompositions of labelled DAGs leading to a recurrence satisfied by the numbers $A_{n,k}$ of DAGs with $n$ vertices including $k$ sources (vertices without any incoming edge).
+He later reformulates those results in terms of a new kind of generating functions, now called graphical generating functions in [36], and also obtains the asymptotic number of size $n$ DAGs.
+Around the same time, Stanley also used a generating function approach in [41] obtained the same results by deriving identities of the chromatic polynomial.
 Robinson also solves the unlabelled case starting from the same ideas but using Burnside’s lemma and cycle index sums to account for the symmetries of these objects.
-He provides a first solution in [Robinson1970] and makes it more computationally tractable in [Robinson1977].
-In the 90’s, Gessel generalised those results, also using the graphical generating function framework in [Gessel1995, Gessel1996] to take into account more parameters and count DAGs by vertices and edges, but also sinks and sources.
+He provides a first solution in [38] and makes it more computationally tractable in [37].
+In the 90’s, Gessel generalised those results, also using the graphical generating function framework in [20, 19] to take into account more parameters and count DAGs by vertices and edges, but also sinks and sources.
 Random sampling
-From the point of view of uniform random generation, the recursive decomposition exhibited by Robinson in [Robinson1973] is interesting as it is amenable to the recursive method pioneered by Nijenhuis and Wilf in [NW1978].
+From the point of view of uniform random generation, the recursive decomposition exhibited by Robinson in [36] is interesting as it is amenable to the recursive method pioneered by Nijenhuis and Wilf in [35].
 This yields a polynomial time algorithm for sampling uniform DAGs with $n$ vertices.
-The analysis of this algorithm has been done in [KM2015] but it had been acknowledged earlier in [MDB2001] although the article proposes an alternative solution.
-Both [KM2015] and [MDB2001] also offer a Markov chain approach to the random sampling problem and an interesting discussion on the pros and cons of both approaches is given in [KM2015].
+The analysis of this algorithm has been done in [31] but it had been acknowledged earlier in [34] although the article proposes an alternative solution.
+Both [31] and [34] also offer a Markov chain approach to the random sampling problem and an interesting discussion on the pros and cons of both approaches is given in [31].
 Remote from the field of combinatorics, the random generation of DAGs is also an active topic in the area of applied statistics and Bayesian inference.
 In this context, DAGs encode a relevant structure in a collection of random variables and the problem of interest is to sample DAGs from a particular distribution related to those random variables.
-To this end, authors resort both to Monte Carlo Markov Chains approaches [KSM2022, KM2017] and methods similar to what is referred to as the recursive method in combinatorics [TVK2020].
-An important point in [KM2017] is better performance can be achieved by using a combination of both approaches, in particular by exploiting the combinatorial properties of DAGs.
-Notable is that sampling from the uniform distribution is tackled as a particular case in [TVK2020] and solved with the asymptotically optimal $O(n^{2})$ complexity at the expense of a $O(n^{3})$ pre-processing step.
+To this end, authors resort both to Monte Carlo Markov Chains approaches [32, 30] and methods similar to what is referred to as the recursive method in combinatorics [43].
+An important point in [30] is better performance can be achieved by using a combination of both approaches, in particular by exploiting the combinatorial properties of DAGs.
+Notable is that sampling from the uniform distribution is tackled as a particular case in [43] and solved with the asymptotically optimal $O(n^{2})$ complexity at the expense of a $O(n^{3})$ pre-processing step.
 Unfortunately, to our knowledge, no efficient uniform random generator of unlabelled has been found yet.
@@
 Indeed, sampling a uniform DAG with a prescribed number of vertices and edges cannot be achieved using the Markov chain approach as it constrains the chain too much, and the formulas of Gessel are not amenable to this either.
-In [KM2015, § 7], the authors provide an interesting discussion on which kind of restrictions can be made on DAGs with the Markov chain approach.
+In [31, § 7], the authors provide an interesting discussion on which kind of restrictions can be made on DAGs with the Markov chain approach.
 They address in particular the case of bounding the number of edges and highlight that the Monte Carlo Markov Chain approach fails when the desired number of edges is too low, thus advocating for having precise combinatorial enumerations.
@@
 This local ordering of the outgoing edges allows to capture more precisely the structure of existing mathematical objects.
-For instance, the compressed formulas and tree-like structures mentioned earlier (see [Ershov1958, Goto1974]) indeed present with an ordering as soon as the underlying tree representation is ordered.
+For instance, the compressed formulas and tree-like structures mentioned earlier (see [13, 22]) indeed present with an ordering as soon as the underlying tree representation is ordered.
 This is the case for most trees used in computer science (e.g.
@@
 As a first step of our analysis, we describe a recursive decomposition scheme that allows us to study DOAGs using tools from enumerative combinatorics.
-This allows us to obtain a recurrence formula for counting them, as well as a polynomial-time uniform random sampler, based on the recursive method from [NW1978], giving full control over their number of vertices and edges.
+This allows us to obtain a recurrence formula for counting them, as well as a polynomial-time uniform random sampler, based on the recursive method from [35], giving full control over their number of vertices and edges.
 Our decomposition is based on a “vertex-by-vertex” approach, that is we remove one vertex at a time and we are able to describe exactly what amount of information is necessary to reconstruct the graph.
@@
 This allows us to transfer our results on DOAGs in the context of labelled DAGs.
-Our new recurrence differs from the formula of Gessel [Gessel1996] in that it does not resort to the inclusion-exclusion principle.
+Our new recurrence differs from the formula of Gessel [19] in that it does not resort to the inclusion-exclusion principle.
 Our approach allows us to obtain an efficient uniform random sampler of labelled DAGs with a prescribed number of vertices, edges, and sources.
@@
 In terms of applications, our random generation algorithms enable to experiment with the properties of the objects they model and with the average complexity of algorithms operating on them.
-A similar approach is for instance taken in [CP2024] where samplers for a realistic class of Git graphs are developed in order to tackle the average complexity of a new algorithm introduced in [Lecoq2024, CDL2022].
-Random testing a also an important application of random sampling, especially as a building block for property-based testing, a now well-established framework pioneered Claessen and Hughes in [CH2000].
-This paper extends an earlier article [GPV2021] with new results on the asymptotics of DOAGs, with an optimal uniform random sampler for the case when the number of edges is not prescribed, and covers a larger class of DOAGs and DAGs by drooping a constraint on the number of sinks.
-For the sake of completeness, the most important results and ideas from [GPV2021] will be recalled in the present paper, but the reader will have to refer the earlier article to get the full proof and algorithmic details.
+A similar approach is for instance taken in [8] where samplers for a realistic class of Git graphs are developed in order to tackle the average complexity of a new algorithm introduced in [33, 7].
+Random testing a also an important application of random sampling, especially as a building block for property-based testing, a now well-established framework pioneered Claessen and Hughes in [5].
+This paper extends an earlier article [18] with new results on the asymptotics of DOAGs, with an optimal uniform random sampler for the case when the number of edges is not prescribed, and covers a larger class of DOAGs and DAGs by drooping a constraint on the number of sinks.
+For the sake of completeness, the most important results and ideas from [18] will be recalled in the present paper, but the reader will have to refer the earlier article to get the full proof and algorithmic details.
 1.3 Outline of the paper
@@
 In Section 3, we quickly go over earlier results regarding the random generation of DOAGs with a prescribed number of vertices, edges, and sources.
-The presentation given in this paper slightly generalises over the algorithm given in [GPV2021] but the ideas and proofs remain unchanged.
+The presentation given in this paper slightly generalises over the algorithm given in [18] but the ideas and proofs remain unchanged.
 Section 4 shows that our approach applies to labelled graphs as well and opens the way for further research regarding this class.
@@
 2 Definition and recursive decomposition
-In this section, we recall a model of directed acyclic graphs called “Directed Ordered Acyclic Graphs” (or DOAGs) that we introduced in [GPV2021].
+In this section, we recall a model of directed acyclic graphs called “Directed Ordered Acyclic Graphs” (or DOAGs) that we introduced in [18].
 It is similar to the classical model of unlabelled DAGs but where, in addition, we have a total order on the outgoing edges of each vertex.
-The presentation we opted for here slightly differs from that of [GPV2021] but essentially defines the same objects, the only difference being that we now allow several sinks for the sake of generality.
+The presentation we opted for here slightly differs from that of [18] but essentially defines the same objects, the only difference being that we now allow several sinks for the sake of generality.
 Definition 1 (Directed Ordered Graph).
@@
   • $E\subset V\times V$ is a set of edges;
-  • for all $v\in V$$\prec_{v}$ is a total order over the set of outgoing edges of $v$;
+  • for all $\in Vv\prec_{v}$ is a total order over the set of outgoing edges of $v$;
   • and $\prec_{\top}$ is a total order over the set of sources of the graph, that is the vertices without any incoming edge.
@@
 We study this class as a whole, however, some sub-classes are also of special interest, in particular for the purpose of modelling compacted data structures.
-Tree structures representing real data, such as XML documents for instance [BLMN2015], are rooted trees.
+Tree structures representing real data, such as XML documents for instance [3], are rooted trees.
 When these trees are compacted, the presence of a root translates into a unique source in the resulting DOAG.
 Similarly, DOAGs with a single sink will arise naturally when compacting trees which bear a single type of leaves.
-In particular the model of compacted binary trees, which can also be seen as a class of cycle-free binary automata, has been shown have unusual combinatorial properties in [EFW2020, EFW2021] and corresponds to a restriction of our model with only binary nodes (and one sink).
+In particular the model of compacted binary trees, which can also be seen as a class of cycle-free binary automata, has been shown have unusual combinatorial properties in [11, 12] and corresponds to a restriction of our model with only binary nodes (and one sink).
 For these reasons, we will also discuss how to approach the sub-classes of DOAGs with a single source and/or a single sink in this document.
@@
 There are 17 of them while there are $95$ DOAGs with $4$ vertices in total.
-Figure 1: All DOAGs with respectly $1$ vertex, $2$ vertices, $3$ vertices, and simultaneously $4$ vertices and $3$ edges.
+Figure 1: All DOAGs with respectly 11 vertex, 22 vertices, 33 vertices, and simultaneously 44 vertices and $3$ edges.
 All edges are implicitly oriented from top to bottom, the blue labels and arrows represent the sources and out-edges orderings (always from left to right).
@@
   • $D^{\prime}$ is a DOAG (obtained by removing $v$ from $D$);
-  • $I$ is any subset of the internal vertices of $D^{\prime}$ (children of $v$ in $D$);
-  • $s$ is any integer between 00 and the number of sources of $D^{\prime}$;
+  • II is any subset of the internal vertices of $D^{\prime}$ (children of $v$ in $D$);
+  • $s$ is any integer between 0 and the number of sources of $D^{\prime}$;
   • and $f:I\to\llbracket 1;s+|I|\rrbracket$ is an injective function (mapping the vertices of $I$ to their positions in the list of children of $v$ in $D$).
@@
 $D_{1,m,k}$ $=\boldsymbol{1}_{\{m=0\,\land\,k=1\}}$ (1)
-$D_{n,m,k}$ =$=\begin{cases}0&\text{when}~{}k\leq 0\\ \sum_{p=0}^{n-k}\sum_{i=0}^{p}D_{n-1,m-p,k-1+p-i}\binom{n-k-p+i}{i}\binom{p}{i}i!&\text{otherwise,}\end{cases}$
+[…]
 where $p=s+i$ corresponds the out-degree of the smallest source.
 The term $\binom{n-k-p+i}{i}=\binom{n-k-s}{i}$ accounts for the choice of the set $I$ and the term $\binom{p}{i}i{!}$ accounts for the number of injective functions $f:I\to\llbracket 1;p\rrbracket$.
-The upper bound on $p$ in the sum is justified by the fact that the out-degree of any vertex can be at most the number of non-sources in the graph, that is $(n-k)$.
-The decomposition scheme presented here differs from the approach described by Robinson in [Robinson1970] as it operates on only one source at a time.
-It is also reminiscent of the peeling processes used in map enumeration where maps are decomposed one face at a time, see for instance [KMSW2019].
+The upper bound on $p$ in the sum is justified by the fact that the out-degree of any vertex can be at most the number of non-sources in the graph, that is (nk)(n-k).
+The decomposition scheme presented here differs from the approach described by Robinson in [38] as it operates on only one source at a time.
+It is also reminiscent of the peeling processes used in map enumeration where maps are decomposed one face at a time, see for instance [27].
 However, the absence of ordering amongst the incoming edges of each vertex in our setup renders those approaches inapplicable as is.
@@
 Note that restricting DOAGs to have only one single sink or one single source ensures that they remain connected, however not all connected DOAGs are obtained this way.
-As another example, DOAGs with out-degree bounded by some constant $d$ are obtained by letting $p$ range from 00 to $\min(n-k,d)$.
+As another example, DOAGs with out-degree bounded by some constant $d$ are obtained by letting $p$ range from 0 to $\min(n-k,d)$.
 The general principle is that the DOAGs whose vertices’ out-degrees are constrained to belong to a given set $\mathcal{P}$, are enumerated by the following generalised recurrence.
 $D^{\mathcal{P}}_{1,m,k}$ $=\boldsymbol{1}_{\{m=0\,\land\,k=1\}}$ (2)
-$D^{\mathcal{P}}_{n,m,k}$ =$=\begin{cases}0&\text{when}~{}k\leq 0\\ \sum_{p\in\mathcal{P}}\sum_{i=0}^{p}D^{\mathcal{P}}_{n-1,m-p,k-1+p-i}\binom{n-k-p+i}{i}\binom{p}{i}i!&\text{otherwise,}\end{cases}$
+[…]
 The first values of the sequence $D_{n,m}=\sum_{k}D_{n,m,k}$ counting DOAGs by number of vertices and edges only are given in Table 1.
@@
 $n$ $D_{n}$ $D_{n,m}=\sum_{k}D_{n,m,k}$ for $m=0,1,2,3,\ldots$
-$1$ $1$ $1$
-$2$ $2$ $1,1$
-$3$ $8$ $1,2,3,2$
-$4$ $95$ $1,3,8,17,27,27,12$
-$5$ $4858$ $1,4,15,48,139,349,718,1136,1272,888,288$
-$6$ $1336729$ $1$, $5$, $24$, $100$, $391$, $1434$, $4868$, $14940$, $40261$, $92493$, $175738$, $266898$, $310096$, $258120$, $136800$, $34560$
+11 11 11
+22 22 $1,1$
+33 88 $1,2,3,2$
+44 $95$ $1,3,8,17,27,27,12$
+55 $4858$ $1,4,15,48,139,349,718,1136,1272,888,288$
+66 $1336729$ 11, 55, $24$, $100$, $391$, $1434$, $4868$, $14940$, $40261$, $92493$, $175738$, $266898$, $310096$, $258120$, $136800$, $34560$
 Table 2: Number of DOAGs with $n$ vertices and a constrained set of allowed degrees.
@@
 (all DOAGs)
-$1$, $2$, $8$, $95$, $4858$, $1336729$, $2307648716$, $28633470321822$, $2891082832793961795$, $2658573971407114263085356$, $24663703371794815015576773905384$, $\ldots$
+11, 22, 88, $95$, $4858$, $1336729$, $2307648716$, $28633470321822$, $2891082832793961795$, $2658573971407114263085356$, $24663703371794815015576773905384$, $\ldots$
 $\mathcal{P}=\mathbb{N}$, $k=1$
 (1 source)
-$1$, $1$, $4$, $57$, $3399$, $1026944$, $1875577035$, $24136664716539$, $2499751751065862022$, $2342183655157963146881571$, $22043872387559770578846044961204$, $\ldots$
+11, 11, 44, $57$, $3399$, $1026944$, $1875577035$, $24136664716539$, $2499751751065862022$, $2342183655157963146881571$, $22043872387559770578846044961204$, $\ldots$
 $\mathcal{P}=\mathbb{N}^{\star}$, $k=1$
 (1 source, 1 sink)
-$1$, $1$, $3$, $37$, $2103$, $627460$, $1142948173$, $14701782996075$, $1522511169925136833$, $1426529804350999351686869$, $13426022673540053054145359653988$, $\ldots$
+11, 11, 33, $37$, $2103$, $627460$, $1142948173$, $14701782996075$, $1522511169925136833$, $1426529804350999351686869$, $13426022673540053054145359653988$, $\ldots$
 $\mathcal{P}=\{0,1,2\}$, $k=1$
 (unary-binary)
-$1$, $1$, $4$, $23$, $191$, $2106$, $29294$, $495475$, $9915483$, $229898277$, $6074257926$, $180460867600$, $5962588299084$, $\ldots$
+11, 11, 44, $23$, $191$, $2106$, $29294$, $495475$, $9915483$, $229898277$, $6074257926$, $180460867600$, $5962588299084$, $\ldots$
 3 Earlier results on counting and recursive sampling
 In this section we summarise our earlier results on the counting and random sampling problem for DOAGs when all three parameters (number of vertices, edges and sources) are fixed.
-The theorems are stated in a slightly more general setting here than in [GPV2021] so as to capture all variants of the model as described in Section 2.1.1.
-However, there is no technical difficulty in the generalisation so that the proofs from [GPV2021] still apply, almost without modification.
-We first present a utility result: computing all the numbers $D_{n,m,k}^{\mathcal{P}}$ up to a certain bound on $n$$m$, and $k$ can be done in polynomial time.
+The theorems are stated in a slightly more general setting here than in [18] so as to capture all variants of the model as described in Section 2.1.1.
+However, there is no technical difficulty in the generalisation so that the proofs from [18] still apply, almost without modification.
+We first present a utility result: computing all the numbers $D_{n,m,k}^{\mathcal{P}}$ up to a certain bound on nnmm, and $k$ can be done in polynomial time.
 This is of moderate interest in itself, but this is a requirement for our samplers, that compute theses values as a pre-processing step.
-Our algorithm is based on the so-called “recursive method” from [NW1978].
+Our algorithm is based on the so-called “recursive method” from [35].
 3.1 Counting
@@
 3.2 Recursive random sampling
-A straightforward application of the recursive method from Nijenhuis and Wilf [NW1978] leads to Algorithm 1, which is presented here in a high level fashion.
+A straightforward application of the recursive method from Nijenhuis and Wilf [35] leads to Algorithm 1, which is presented here in a high level fashion.
 Algorithm 1 Recursive uniform sampler of DOAGs
@@
 6:   pick $(p,i)$ with probability ${D_{n-1,m-p,k-1+p-i}^{\mathcal{P}}\binom{n-k-p+i}{i}\binom{p}{i}i{!}}/{D_{n,m,k}^{\mathcal{P}}}$
-[…]
+7:   $D^{\prime}\leftarrow\text{{UnifDOAG${}^{\mathcal{P}}$}($n-1,m-p,k-1+p-i$)}$
 8:   $I\leftarrow$ a uniform subset of size $i$ of the inner vertices of $D^{\prime}$
@@
 10:   return $\operatorname{\mathrm{decomp}}^{-1}(D^{\prime},p-i,I,f)$
-In [GPV2021, §3], we discuss how to implement Algorithm 1 efficiently.
+In [18, §3], we discuss how to implement Algorithm 1 efficiently.
 In particular we suggest a data-structure to represent DOAGs that allows for an efficient selection of the subset $I$ at line 8 and the function $f$ at line 9.
@@
 As mentioned above, the numbers $D_{n,m,k}^{\mathcal{P}}$ either have to be pre-computed for Algorithm 1 to work, or must be lazily computed and memoised on the fly.
-In practice, pre-computing all the necessary numbers to sample a uniform DOAG with $50$ vertices (without any constraint on $m$ and on the out-degree) using our library already takes about $8$ seconds on a standard laptop.
+In practice, pre-computing all the necessary numbers to sample a uniform DOAG with $n=50$ vertices (without any constraint on $m$ and on the out-degree) using our library already takes about $8$ seconds on a standard laptop.
 This running time rapidly increases, which makes the cost generating large structures prohibitive.
@@
 Also note that in order for the algorithm to be made generic in $\mathcal{P}$, we only have to use the sequence $D_{n,m,k}^{\mathcal{P}}$ rather than $D_{n,m,k}$, which reflects the generality of the recursive method.
-Four random DOAGs, drawn using Algorithm 1 with $\mathcal{P}=\{0,1,2\}$$n=1250$$m=1300$ and $k=1$ are pictured in Figure 3(b).
+Four random DOAGs, drawn using Algorithm 1 with $\mathcal{P}=\{0,1,2\}$n=1250n=1250m=1300m=1300 and $k=1$ are pictured in Figure 3(b).
 As a comparison, a truncated version of the Git history of the linux kernel is pictured on the left in Figure 3(a).
@@
 In this section we demonstrate how our decomposition scheme can be applied to the classical model of labelled DAGs to obtain new recurrences on known sequences.
-In the 1990s, Gessel [Gessel1996] already obtained equations allowing to count labelled DAGs by vertices, edges, and sources (and also sinks actually) using a generating functions approach.
+In the 1990s, Gessel [19] already obtained equations allowing to count labelled DAGs by vertices, edges, and sources (and also sinks actually) using a generating functions approach.
 These equations involve the inclusion-exclusion principle which has one drawback: they are usually not amenable to efficient random generation.
@@
 As a consequence, we can obtain an efficient random sampler of DAGs with full control over the number of vertices, sources, and edges.
-As in Section 2, we present here a slight generalisation of the formula given in [GPV2021, §4] allowing to capture various classes of labelled DAGs.
-We omit the proofs here as they follow a straightforward adaptation of the arguments given in [GPV2021] and the recursive method.
+As in Section 2, we present here a slight generalisation of the formula given in [18, §4] allowing to capture various classes of labelled DAGs.
+We omit the proofs here as they follow a straightforward adaptation of the arguments given in [18] and the recursive method.
 4.1 Recursive decomposition
@@
 Let $D$ denote one such DAG and let $v$ denote its distinguished source.
-Removing the distinguished source in $D$ and decrementing the labels of the vertices of higher label than $v$ by one yields a regular vertex-labelled DAG $D^{\prime}$ with $n-1$ vertices.
+Removing the distinguished source in $D$ and decrementing the labels of the vertices of higher label than $v$ by one yields a regular vertex-labelled DAG $D^{\prime}$ with n1n-1 vertices.
 Moreover, the three pieces of information that are necessary to reconstruct the source are the following:
@@
 $A_{1,m,k}^{\mathcal{P}}$ $=\boldsymbol{1}_{\{m=0~\wedge~k=1\}}$ (3)
-$kA_{n,m,k}^{\mathcal{P}}$ =$=\begin{cases}n\sum_{p\in\mathcal{P}\cap\llbracket 0;n-k\rrbracket}\sum_{i=0}^{p}A_{n-1,m-p,k-1+p-i}^{\mathcal{P}}\binom{n-k-p+i}{i}\binom{k-1+p-i}{p-i}&\text{if~{}$1\leq k$}\\ 0&\text{otherwise.}\end{cases}$
+[…]
 In the last equation:
@@
   • and the two binomial coefficient count the number of ways to select the subsets $I$ and $S$.
-When $\mathcal{P}=\mathbb{N}$, we recover the sequence counting all labelled DAGs, known as A003024 in the OEIS and first enumerated in [Robinson1970, Stanley1973, Robinson1973].
-For $\mathcal{P}=\mathbb{N}^{\star}$ and with $k=1$, we find the number of labelled DAGs with a single source and a single sink, known to Gessel in [Gessel1995, Gessel1996] and stored at A165950 in the OEIS.
+When $\mathcal{P}=\mathbb{N}$, we recover the sequence counting all labelled DAGs, known as A003024 in the OEIS and first enumerated in [38, 41, 36].
+For $\mathcal{P}=\mathbb{N}^{\star}$ and with $k=1$, we find the number of labelled DAGs with a single source and a single sink, known to Gessel in [20, 19] and stored at A165950 in the OEIS.
 4.2 Random generation
@@
    pick $(p,i)$ with probability $\frac{A_{n-1,m-p,k-1+p-i}^{\mathcal{P}}\binom{n-k-p+i}{i}\binom{k-1+p-i}{p-i}}{A_{n,m,k}^{\mathcal{P}}}$
-         $D^{\prime}$leftarrow unifDAGP($n-1,m-p,k-1+p-i$)
-         $I$leftarrow a uniform subset of size $i$ of the inner vertices of $D^{\prime}$
-         $S$leftarrow a uniform subset of size $(p-i)$ the sources of $D^{\prime}$
-         $\ell$leftarrow Unif($\llbracket 1;n\rrbracket$)
+   $D^{\prime}$ $\leftarrow$ unifDAGP(n1,mp,k1+pin-1,m-p,k-1+p-i)
+   II $\leftarrow$ a uniform subset of size $i$ of the inner vertices of $D^{\prime}$
+   $S$ $\leftarrow$ a uniform subset of size (pi)(p-i) the sources of $D^{\prime}$
+   $\ell$ $\leftarrow$ Unif($\llbracket 1;n\rrbracket$)
    relabel $D^{\prime}$ by adding one to all labels $\ell^{\prime}\geq\ell$
@@
 We then consider its transition matrix using these labels as indices.
-Usually, the transition matrix of a directed graph $D$ is defined as the matrix ${(a_{i,j})}_{1\leq i,j\leq n}$ such that $a_{i,j}$ is $1$ if there is an edge from vertex $i$ to vertex $j$ in $D$, and 00 otherwise.
+Usually, the transition matrix of a directed graph $D$ is defined as the matrix ${(a_{i,j})}_{1\leq i,j\leq n}$ such that $a_{i,j}$ is $1$ if there is an edge from vertex $i$ to vertex $j$ in $D$, and 0 otherwise.
 This representation encodes the set of the edges of a DAG but not the edge ordering of DOAGs.
@@
 The size of a variation is its length.
-[…]
+For instance, the sequence $(6,2,3,0,0,1,4,0,5)$ is a variation of size $9$ but the sequences $(1,0,3)$ and $(1,0,3,2,3)$ are not variations because the number $2$ is missing in the first one and the second contains two occurrences of the number $3$.
+Variations can also be defined as interleavings of a permutation with a sequence of zeros.
+One of the earliest references to these objects dates back to 1659 in Izquierdo’s Pharus Scientiarum [25, Disputatio 29].
+They also appear in Stanley’s book as the second entry of his Twelvefold Way [42, page 79], a collection of twelve basic but fundamental counting problems.
+Knuth gives a few ancient references on this topic in [29] and in an quote (without reference) that can be found on the OEIS page of variations at A007526.
+Variations are relevant to our problem as they naturally appear as rows of the labelled transition matrices defined in this section.
+Some of their combinatorial properties will be exhibited in the next section.
 Definition 5 (Variation matrix).
@@
   • it is strictly upper triangular;
-  • for all $11\leq i\leq n-$ 1, the sub-row ${(a_{i,j})}_{i<j\leq n}$ is a variation (of size $n-i$).
-From an enumerative point of view, a variation matrix can be seen as a sequence of variations $(v_{1},v_{2},\ldots,v_{n-1})$ where for all $11\leq i\leq n-$ 1, the variation $v_{i}$ has size $i$.
+  • for all $1\leq i\leq n-1$, the sub-row ${(a_{i,j})}_{i<j\leq n}$ is a variation (of size nin-i).
+From an enumerative point of view, a variation matrix can be seen as a sequence of variations $(v_{1},v_{2},\ldots,v_{n-1})$ where for all $1\leq i\leq n-1$, the variation $v_{i}$ has size $i$.
 We have established that all labelled transition matrices of DOAGs are variation matrices.
@@
 Another important remark is that, when several underlined cells occur on the same row $i$ in the matrix, they correspond to several sources that are discovered at the same decomposition step of the DOAG (upon removing the same source).
-Recall that the decomposition
 Figure 7: Same-row underlined entries are sorted.
-algorithm sorts the labels of these new sources by following the total order of the outgoing edges of vertex $i$.
+Recall that the decomposition algorithm sorts the labels of these new sources by following the total order of the outgoing edges of vertex $i$.
 This implies that the underlined entries within the same row have to be increasingly sorted (from left to right).
@@
 All labelled transition matrices of DOAGs are variation matrices.
-Furthermore, let $A={(a_{i,j})}_{1\leq i,j\leq n}$ be a variation matrix, and for all $j\in\llbracket 1;n\rrbracket$, let $b_{j}$ denote the largest $i\leq n$ such that $a_{i,j}>0$ if such an index exists and 00 otherwise.
+Furthermore, let $A={(a_{i,j})}_{1\leq i,j\leq n}$ be a variation matrix, and for all $j\in\llbracket 1;n\rrbracket$, let $b_{j}$ denote the largest $i\leq n$ such that $a_{i,j}>0$ if such an index exists and 0 otherwise.
 Then, $A$ is the labelled transition matrix of some DOAG if and only if the two following properties hold:
@@
 There exists a constant $c>0$ such that the number $D_{n}$ of DOAGs with $n$ vertices and the number $D_{n}^{\star}$ of such DOAGs having only one source satisfy
-[…]
+$D_{n}$ $=\left(1+O\!\left(n^{-1}\right)\right)D_{n}^{\star}=\frac{c}{\sqrt{n}}\ e^{n-1}\ \text{¡}{(n-1)}!\left(1+O\!\left(n^{-1}\right)\right)$
 $\sim c\cdot e^{\zeta^{\prime}(-1)-1}\cdot n^{-7/12}\cdot{\big(e\sqrt{2\pi}\big)}^{n}\!\cdot e^{-\frac{3}{4}n^{2}}\cdot n^{n^{2}/2}$
-[…]
+where $\text{¡}{k}!=\prod_{i=0}^{k}{i!}$ denotes the super factorial of $k$.
 The super factorial provides a concise way to express this equivalent and also reflects the relation between DOAGs and variation matrices, which will be further developed in this section.
@@
 For all $n\geq 1$, the number $D_{n}$ of DOAGs of size $n$ satisfies
-[…]
-[…]
-[…]
-Lemma 2 (Asymptotic estimation of the super-factorial [Barnes1900]).
+$D_{n}\leq\text{¡}{(n-1)}!e^{n-1}$
+where $\text{¡}{k}!=\prod_{i=0}^{k}{i!}$ denotes the super factorial of $k$.
+The term “super factorial” seems to have been coined by Sloane and Plouffe in [39, page 228] but this sequence had been studied before that, in 1900, by Barnes [2] as the integer values of the “G-function”.
+In fact, if $G(z)$ denotes the complex-valued G-function of Barnes, we have the identity $G(n+2)=\text{¡}{n}!$ for all integer $n$.
+Barnes also gives the following equivalent.
+Lemma 2 (Asymptotic estimation of the super-factorial [2]).
 When $n\to\infty$, we have
-[…]
+$\text{¡}{(n-1)}!=G(n+1)\sim e^{\zeta^{\prime}(-1)}\cdot n^{-1/12}\cdot{\big(\sqrt{2\pi}\big)}^{n}\cdot e^{-\frac{3}{4}n^{2}}\cdot n^{n^{2}/2}$
 where $\zeta$ denotes the Riemann zeta function.
@@
 Proof.
-Let $0\leq p\leq n$, a variation of size $n$ containing exactly $p$ zeros is the interleaving of a permutation of size $(n-p)$ with an array of zeros of size $p$. As a consequence
+Let $0\leq p\leq n$, a variation of size $n$ containing exactly $p$ zeros is the interleaving of a permutation of size (np)(n-p) with an array of zeros of size $p$. As a consequence
 $v_{n,p}=\binom{n}{p}(n-p)!=\frac{n!}{p!}.$
@@
 We thus have the following upper bound for $D_{n}$:
-[…]
+$D_{n}\leq\prod_{i=1}^{n-1}v_{n-i}\leq\prod_{i=1}^{n-1}e\cdot(n-i)!=\text{¡}{(n-1)}!e^{n-1}.\qed$
 Figure 9: Lower-bound on the set of labelled transition matrices.
@@
 There exists a constant $A>0$ such that for all $n\geq 1$, we have
-[…]
+$\frac{A}{n}\text{¡}{(n-1)}!e^{n-1}\leq D_{n}.$
 Proof.
-In a labelled transition matrix with positive values on the super-diagonal, the $i$-th row can be seen as a variation of size $(n-i)$ that does not start with a zero.
-Moreover, the number of variations of size $n$ starting with a zero is actually the number of variations of size $(n-1)$ so that the number of possibilities for the $i$-th row of the matrix we count here is $(v_{n-i}-v_{n-i-1})$.
+In a labelled transition matrix with positive values on the super-diagonal, the $i$-th row can be seen as a variation of size (ni)(n-i) that does not start with a zero.
+Moreover, the number of variations of size $n$ starting with a zero is actually the number of variations of size (n1)(n-1) so that the number of possibilities for the $i$-th row of the matrix we count here is $(v_{n-i}-v_{n-i-1})$.
 In addition, by Lemma 3, we also have that
 $v_{n}-v_{n-1}=e\cdot n!-e\cdot(n-1)!+o(1)=e\cdot n!\cdot\frac{n-1}{n}\left(1+O\left(\frac{1}{n!}\right)\right).$ (4)
-Note that when $i=n-1$, we have $v_{n-i}-v_{n-i-1}=v_{1}-v_{0}=1$.
-Indeed, the row of index $i=n-1$ contains only the number $1$ in the super diagonal since at the last step of the decomposition, we have two connected vertices and there is only one such DOAG.
+Note that when i=n1i=n-1, we have $v_{n-i}-v_{n-i-1}=v_{1}-v_{0}=1$.
+Indeed, the row of index i=n1i=n-1 contains only the number $1$ in the super diagonal since at the last step of the decomposition, we have two connected vertices and there is only one such DOAG.
 Setting aside this special case, which does not contribute to the product, we get the following lower bound for $D_{n}$:
-[…]
+$D_{n}\geq\prod_{i=1}^{n-2}(v_{n-i}-v_{n-i-1})=e^{n-2}\text{¡}{(n-1)}!\prod_{i=1}^{n-2}\frac{n-1-i}{n-i}\prod_{i=1}^{n-1}\left(1+O\left(\frac{1}{(n-i)!}\right)\right)$ (5)
 where the first product telescopes and yields $\frac{1}{n-1}$ and the second one converges to a constant as $n\to\infty$. This allows to conclude the proof the lemma. ∎
-[…]
+Although they are not precise enough to obtain an asymptotic equivalent for the sequence $D_{n}$, these two bounds already give us a good understanding of the behaviour of $D_{n}$.
+First of all, they let appear a “dominant” term of the form $\text{¡}{(n-1)}!e^{n-1}$, which is uncommon in combinatorial enumeration.
+And second, it tells us we only make a relative error of the order of $O(n)$ when approximating $D_{n}$ by $\text{¡}{(n-1)}!\cdot e^{n-1}$.
+We will prove an asymptotic equivalent for the remaining polynomial term, but in order to obtain this, we first need to slightly refine our lower bound so that the “interval” between our two bounds is a little narrower than $O\!\left(n\right)$.
 Lemma 5 (A better lower bound for the number of DOAGs).
 There exists a constant $B>0$ such that, for all $n\geq 1$, we have
-[…]
+$D_{n}\geq B\frac{\ln(n)}{n}\text{¡}{(n-1)}!e^{n-1}.$
 Proof.
 In order to obtain this lower bound, we count the number of valid labelled transition matrices such that all but exactly one of the cells on the super-diagonal have non-zero values.
-Furthermore, in order to avoid having to deal with border cases, we assume that the unique zero value on the super-diagonal appears between $i=2$ and $i=n-2$.
-Let $22\leq i\leq n-$ 2 and let us consider those matrices ${(a_{p,q})}_{1\leq p,q\leq n}$ such that $a_{i,i+1}=0$.
+Furthermore, in order to avoid having to deal with border cases, we assume that the unique zero value on the super-diagonal appears between $i=2$ and i=n2i=n-2.
+Let $2\leq i\leq n-2$ and let us consider those matrices ${(a_{p,q})}_{1\leq p,q\leq n}$ such that $a_{i,i+1}=0$.
 Those matrices are illustrated in Figure 10
@@
   1. 1.
-    On row $i-1$, the two first cells on the right of the diagonal ($a_{i-1,i}$ and $a_{i-1,i+1}$) must have positive values and must be in increasing order.
+    On row i1i-1, the two first cells on the right of the diagonal ($a_{i-1,i}$ and $a_{i-1,i+1}$) must have positive values and must be in increasing order.
 In the case of $a_{i-1,i}$, this is because it is on the super diagonal.
@@
 Otherwise the condition from Theorem 3 are violated.
-In terms of DOAGs, this means that the vertex $i-1$ produces two new sources when it is removed but vertex $i$ produces none.
+In terms of DOAGs, this means that the vertex i1i-1 produces two new sources when it is removed but vertex $i$ produces none.
   2. 2.
-    On row $i$, any variations of size $(n-i)$ starting by a zero is allowed.
+    On row $i$, any variations of size (ni)(n-i) starting by a zero is allowed.
 We get the number of variations of size $n$ starting by two increasing positive values (condition 1 above) by inclusion-exclusion. That is,
@@
 $=\frac{(v_{n-i-1}-2v_{n-2-i}+v_{n-3-i})v_{n-i-1}}{2(v_{n-i+1}-v_{n-i})(v_{n-i}-v_{n-i-1})}\cdot\prod_{p=1}^{n-1}(v_{n-p}-v_{n-p-1}).$
-By summing over $22\leq i\leq n-$ 2, we get
+By summing over $2\leq i\leq n-2$, we get
 $\sum_{i=2}^{n-2}\frac{(v_{n-i-1}-2v_{n-2-i}+v_{n-3-i})v_{n-i-1}}{2(v_{n-i+1}-v_{n-i})(v_{n-i}-v_{n-i-1})}\cdot\prod_{p=1}^{n-1}(v_{n-p}-v_{n-p-1}).$ (6)
@@
 $\sum_{i=2}^{n-2}\frac{(v_{i-1}-2v_{i-2}+v_{i-3})v_{i-1}}{2(v_{i+1}-v_{i})(v_{i}-v_{i-1})}\sim\sum_{i=2}^{n-2}\frac{1}{2i}\sim\frac{\ln(n)}{2}$
-[…]
+In addition, we know from the proof of Lemma 4 that the product in equation (6) is equivalent to $\frac{c}{n}e^{n-1}\text{¡}{(n-1)}!$ for some constant $c$, which allows to conclude. ∎
 6.2 Obtaining the polynomial term by bootstrapping
 Let us denote $P_{n}$ the polynomial term in $D_{n}$, that is the quantity
-[…]
+$P_{n}\overset{\text{def}}{=}\frac{D_{n}}{\text{¡}{(n-1)}!e^{n-1}}.$
 We have proved above that for some constant $B>0$, we have $B\frac{\ln(n)}{n}\leq P_{n}\leq 1$. A consequence of these inequalities is that for all $k\in\mathbb{Z}$, we have
@@
 In this case, the first source is not connected to the second vertex and the matrix has thus more than one source.
-The first row of such a matrix is a variation of size $(n-2)$ and the lower part ${(a_{i,j})}_{2\leq i,j\leq n}$ encodes a DOAG of size $(n-1)$, the DOAG obtained by removing the first source.
-However, it is important to note that not all combinations of a size-$(n-2)$ variation and a size-$(n-1)$ matrix yield a valid size-$n$ labelled transition matrix.
+The first row of such a matrix is a variation of size (n2)(n-2) and the lower part ${(a_{i,j})}_{2\leq i,j\leq n}$ encodes a DOAG of size (n1)(n-1), the DOAG obtained by removing the first source.
+However, it is important to note that not all combinations of a size-(n2)(n-2) variation and a size-(n1)(n-1) matrix yield a valid size-$n$ labelled transition matrix.
 For instance, a variation of the form $v=(0,1,0,2,\ldots)$ and a lower matrix with at least three sources cannot be obtained together as they would violate the constraints of Theorem 3.
 Case 2: $\boldsymbol{a_{1,2}>0\wedge a_{2,3}>0}$.
-In this case, the first row is a variation of size $(n-1)$ starting by a positive value, and the lower part ${(a_{i,j})}_{2\leq i,j\leq n}$ encodes a DOAG of size $(n-1)$ with exactly one source, again obtained by removing the first source.
+In this case, the first row is a variation of size (n1)(n-1) starting by a positive value, and the lower part ${(a_{i,j})}_{2\leq i,j\leq n}$ encodes a DOAG of size (n1)(n-1) with exactly one source, again obtained by removing the first source.
 This second fact is a direct consequence of $a_{2,3}>0$.
@@
 Case 3: $\boldsymbol{a_{1,2}>0\wedge a_{2,3}=0\wedge a_{3,4}>0}$.
-In this case the lower part ${(a_{i,j})}_{3\leq i,j\leq n}$ encodes a DOAG of size $n-2$ with exactly one source, the first row is necessarily a variation of size $(n-1)$, starting with two positive increasing values, and the second row is a variation of size $(n-2)$ starting by a zero.
+In this case the lower part ${(a_{i,j})}_{3\leq i,j\leq n}$ encodes a DOAG of size n2n-2 with exactly one source, the first row is necessarily a variation of size (n1)(n-1), starting with two positive increasing values, and the second row is a variation of size (n2)(n-2) starting by a zero.
 Here again this decomposition is exact: all such triplets can be obtained here.
@@
 Finally, this case captures all the remaining matrices.
-The first row is a variation of size $(n-1)$, the second and third rows are variations of sizes $(n-2)$ and $(n-3)$ starting with a zero, and the lower part ${(a_{i,j})}_{4\leq i,j\leq n}$ encodes a size-$(n-3)$ DOAG.
+The first row is a variation of size (n1)(n-1), the second and third rows are variations of sizes (n2)(n-2) and (n3)(n-3) starting with a zero, and the lower part ${(a_{i,j})}_{4\leq i,j\leq n}$ encodes a size-(n3)(n-3) DOAG.
 Of course, not all such quadruples can be obtained, but this over-approximation will be enough for our proof.
@@
 The symbols $\mathcal{D}$ and $\mathcal{D}$$^{\star}$ respectively represent the set of all possible DOAG labelled transition matrices the set of all of those matrices such that $a_{1,2}>0$.
-The stars ()($\star$)( ⋆ ) represent strictly positive values.
+The stars $(\star)$ represent strictly positive values.
 We compute the contributions to $D_{n}$ coming from each of these four terms described above.
@@
 $D_{n}=D_{n}^{\star}+O(v_{n-2}D_{n-1}).$ (8)
-Note that the big-$O$ term comes from the fact that not all pairs made of a size-$(n-2)$ variation and a size-$(n-1)$ labelled transition matrix can be obtained this way, as discussed in the first case above.
+Note that the big-$O$ term comes from the fact that not all pairs made of a size-(n2)(n-2) variation and a size-(n1)(n-1) labelled transition matrix can be obtained this way, as discussed in the first case above.
 We could actually have written $0\leq D_{n}-D_{n}^{\star}\leq v_{n-2}D_{n-1}$.
@@
 $D_{n}^{\star}=(v_{n-1}-v_{n-2})D_{n-1}^{\star}+\frac{v_{n-1}-2v_{n-2}+v_{n-3}}{2}v_{n-3}D_{n-2}^{\star}+O\!\left(v_{n-1}v_{n-3}v_{n-4}D_{n-3}\right).$ (9)
-[…]
+Let us introduce the polynomial term $P_{n}^{\star}$ of $D_{n}^{\star}$ defined by $P_{n}^{\star}=D_{n}^{\star}/e^{n-1}\text{¡}{(n-1)}!$. By normalising equation (8) and using equation (7) we have
 $P_{n}^{\star}=P_{n}+O\!\left(\frac{v_{n-2}}{e(n-1)!}P_{n-1}\right)=P_{n}+O\!\left(\frac{P_{n-1}}{n}\right)=P_{n}+o\!\left(P_{n}\right).$
-[…]
+In other words, we have that $P_{n}$ and $P_{n}^{\star}$ are equivalent. Then, by normalising equation (9) by $e^{n-1}\text{¡}{(n-1)}!$, we obtain the following asymptotic expansion
 $P_{n}^{\star}=\left(1-\frac{1}{n}+O\!\left(\frac{1}{n^{2}}\right)\right)P_{n-1}^{\star}+\frac{1}{2n}\left(1+O\!\left(\frac{1}{n}\right)\right)P_{n-2}^{\star}+O\!\left(\frac{P_{n-3}}{n^{2}}\right)\cdot$ (10)
@@
 This is expected because we eliminated the parameter $m$.
-[…]
+Note that the $D_{n}^{\star}$ sequence from Theorem 4 corresponds to $D_{n,1}$ and that $D_{n}=\sum_{k=1}^{n}\hskip-1.0ptD_{n,k}$.
+Using the numbers computed by this algorithm, we plotted the first $250$ values of the sequences $D_{n}$ and $D_{n}^{\star}$ normalised by $n^{-1/2}e^{n-1}\text{¡}{(n-1)}!$ which shows the convergence to the constant $c$ from Theorem 4.
+We also note that the convergence looks faster for the sequence $D_{n}^{\star}$.
+This suggests that the constant can be approximated by $c\approx 0.4967$.
+Figure 12 shows this plot as well as a zoomed-in version near $\frac{1}{2}$ for $n\geq 200$.
 05010015020025000.20.40.60.81
 2002102202302402500.490.4950.50.5050.51
-[…]
+Figure 12: The first values of the sequences ${\frac{D_{n}\sqrt{n}}{\text{¡}{(n-1)}!e^{n-1}}}$ in blue and ${\frac{D_{n}^{\star}\sqrt{n}}{\text{¡}{(n-1)}!e^{n-1}}}$ in red.
 6.4 Asymptotic behaviour of some parameters
@@
 In this case the sub-matrix ${(a_{i,j})}_{k\leq i,j\leq n}$ can be re-interpreted as a DOAG with only one source.
-Indeed, the condition $a_{k,k+1}$ means that upon removing the $(k-1)$ first sources of the DOAG, the decomposition algorithm does not produce any new source, leaving us with a single-source DOAG.
-We can characterise those matrices: they are made of $(k-1)$ variations of size $(n-k)$ in the first rows, and a size-$(n-k+1)$ labelled transition matrix corresponding to a single source DOAG below them.
-Any combination of $(k-1)$ such variations and such a matrix can be obtained.
+Indeed, the condition $a_{k,k+1}$ means that upon removing the (k1)(k-1) first sources of the DOAG, the decomposition algorithm does not produce any new source, leaving us with a single-source DOAG.
+We can characterise those matrices: they are made of (k1)(k-1) variations of size (nk)(n-k) in the first rows, and a size-(nk+1)(n-k+1) labelled transition matrix corresponding to a single source DOAG below them.
+Any combination of (k1)(k-1) such variations and such a matrix can be obtained.
 Case $\boldsymbol{a_{k,k+1}=0}$.
 On the other hand we have the matrices such that $a_{k,k+1}=0$.
-In this case, the $(k-1)$ first rows of the matrix are still variations of size $(n-k)$.
+In this case, the (k1)(k-1) first rows of the matrix are still variations of size (nk)(n-k).
 The lower part ${(a_{i,j})}_{k\leq i,j\leq n}$ can be seen as a DOAG with at least two sources because its first two columns are empty.
-Note that here, depending of the $(k-1)$ top variations, we may have restriction on which DOAGs may appear in the lower part.
-For instance, if $k=2$ and the first row is $(0,0,1,0,0,\ldots,0)$, then the DOAG of size $(n-1)$ without any edge cannot appear in the lower part.
+Note that here, depending of the (k1)(k-1) top variations, we may have restriction on which DOAGs may appear in the lower part.
+For instance, if $k=2$ and the first row is $(0,0,1,0,0,\ldots,0)$, then the DOAG of size (n1)(n-1) without any edge cannot appear in the lower part.
 This dichotomy is pictured in Figure 13.
@@
 $\frac{1}{v_{n}}\sum_{p=0}^{n}pv_{n,p}=\frac{n!}{v_{n}}\sum_{p=0}^{n}\frac{p}{p!}=\frac{n!}{v_{n}}\sum_{p=1}^{n}\frac{1}{(p-1)!}\underset{n\to\infty}{\to}1$
-where $v_{n,p}$ is the number of variations of size $n$ having exactly $p$ zeros, which is equal to $n!/p{!}$ by Lemma 3.
+where $v_{n,p}$ is the number of variations of size $n$ having exactly $p$ zeros, which is equal to n!/p!n!/p{!} by Lemma 3.
 Moreover, the tail of their probability distribution is more than exponentially small:
@@
 For all integer $p\geq 0$, by inclusion, we have that the number of DOAG labelled transition matrices with exactly $p$ zeros strictly above the diagonal is upper-bounded by the number $\operatorname{\mathrm{VM}}_{n,p}$ of variation matrices with the same property.
-Moreover, given a size-$(n-1)$ vector $(p_{1},p_{2},\ldots,p_{n-1})$ of non-negative integers such that for all $i$$p_{i}\leq i$, the number of such variation matrices with exactly $p_{n-i}$ zeros in the $i$-th line is
-[…]
+Moreover, given a vector $(p_{1},p_{2},\ldots,p_{n-1})$ of non-negative integers such that for all $ip_{i}\leq i$, the number of such variation matrices with exactly $p_{n-i}$ zeros in the $i$-th line is
+$\prod_{i=1}^{n-1}v_{i,p_{i}}=\prod_{i=1}^{n-1}\frac{i!}{p_{i}!}=\text{¡}{(n-1)}!\prod_{i=1}^{n-1}\frac{1}{p_{i}!}.$
 By summation over all such vectors such that $\sum_{i=1}^{n-1}p_{i}=p$, we get an expression for $\operatorname{\mathrm{VM}}_{n,p}$:
-[…]
+$\operatorname{\mathrm{VM}}_{n,p}=\text{¡}{(n-1)}!\sum_{\begin{subarray}{c}p_{1}+p_{2}+\cdots+p_{n-1}=p\\ \text{for all~$i$},0\leq p_{i}\leq i\end{subarray}}\prod_{i=1}^{n-1}\frac{1}{p_{i}!}\leq\text{¡}{(n-1)}!\sum_{\begin{subarray}{c}p_{1}+p_{2}+\cdots+p_{n-1}=p\\ \text{for all~$i$},0\leq p_{i}\end{subarray}}\prod_{i=1}^{n-1}\frac{1}{p_{i}!}\cdot$
 In the first sum we have the constraint $p_{i}\leq i$ because a variation has at most $i$ zeros.
@@
 It follows that
-[…]
+$\operatorname{\mathrm{VM}}_{n,p}\leq\text{¡}{(n-1)}!\frac{{(n-1)}^{p}}{p!}\cdot$
 As a consequence, we have the following bound for the probability that a uniform DOAG of size $n$ has at most $\binom{n}{2}-q$ zeros:
-[…]
-The sum in the last equation is the remainder in the Taylor expansion of order $(q-1)$ of the function $e^{x}$ near zero, evaluated at the point $(n-1)$.
+$\mathbb{P}_{n}[\text{a uniform DOAG has at most~$\binom{n}{2}-q$ zeros}]\leq\frac{\text{¡}{(n-1)}!}{D_{n}}\sum_{p\geq q}\frac{{(n-1)}^{p}}{p!}\cdot$ (15)
+The sum in the last equation is the remainder in the Taylor expansion of order (q1)(q-1) of the function $e^{x}$ near zero, evaluated at the point (n1)(n-1).
 By using the integral form of this remainder, we have that
@@
 Finally, by using this estimate inside equation (15), and by using Theorem 4 for estimating $D_{n}$, we get that there exists a constant $c^{\prime}>0$ such that
-[…]
+$\mathbb{P}_{n}[\text{a uniform DOAG has at most~$\binom{n}{2}-\lambda(n-1)$ zeros}]$ $\leq\frac{\text{¡}{(n-1)}!e^{n-1}}{D_{n}}\sum_{p\geq q}\frac{{(n-1)}^{q}}{q!}$
 $\leq\frac{c^{\prime}}{\sqrt{\lambda}}{\left(\frac{e^{\lambda}}{\lambda^{\lambda}}\right)}^{\mathclap{\,\,n-1}}\cdot$
@@
 Indeed,
-$\mathbb{P}[\text{a uniform variation of size~{}$n$ has~{}$p$ zeros}]=\frac{v_{n,p}}{v_{n}}\propto\frac{\boldsymbol{1}_{\{0\leq p\leq n\}}}{p!}$
+[…]
 by Lemma 3.
-A possible way to generate a uniform variation is thus to draw a Poisson variable $p$ of parameter $1$ conditioned to be at most $n$, and then shuffling a size $p$ array of zeros concatenated with the identity permutation using the Fisher-Yates algorithm [FY1948].
+A possible way to generate a uniform variation is thus to draw a Poisson variable $p$ of parameter $1$ conditioned to be at most $n$, and then shuffling a size $p$ array of zeros concatenated with the identity permutation using the Fisher-Yates algorithm [14].
 This is described in Algorithm 3.
@@
 3:function UnifVariation($n$)
-4:     $p$leftarrow BoundedPoisson($1,n$)
-5:     $A$leftarrow $[0,0,\ldots,0,1,2,\ldots,n-p]$$\triangleright$ array of length $n$, starting with $p$ zeros
-6:     for $i=0$ to $n-2$ do
-7:         $r$leftarrow Unif($\llbracket i;n-1\rrbracket$)
-8:         $A[r]\leftrightarrow A[i]$triangleright Swap entries of indices $r$ and $i$
-9:     return $A$
+4:  $p$ $\leftarrow$ BoundedPoisson($1,n$)
+5:  AA $\leftarrow$ $0,0,\ldots,0,1,2,\ldots,n-p]\triangleright$ array of length $n$, starting with $p$ zeros
+6:  for $i=0$ to n2n-2 do
+7:   $r$ $\leftarrow$ Unif($\llbracket i;n-1\rrbracket$)
+8:   $A[r]\leftrightarrow A[i]$ $\triangleright$ Swap entries of indices $r$ and $i$
+9:  return AA
 Regarding the generation of the bounded Poisson variable (performed at line 4), an efficient approach is to generate regular (unbounded) Poisson variables until a value less than $n$ is found.
@@
 As a consequence, the expected number of tries of a rejection procedure for sampling conditioned $\text{Poisson}(1)$ variables is $\frac{1}{p_{n}}\leq\frac{e}{2}$.
-The algorithm described by Knuth in [Knuth1997, page 137] is suitable for our use-case since our Poisson parameter ($1$ here) is small.
+The algorithm described by Knuth in [28, page 137] is suitable for our use-case since our Poisson parameter ($1$ here) is small.
 Furthermore it can be adapted to stop early when values strictly larger than $n$ are found.
@@
 4:  repeat
-5:         $k$leftarrow 00
-6:         $p$leftarrow Unif($[0;1]$)
+5:   $k$ $\leftarrow$ 0
+6:   $p$ $\leftarrow$ Unif($[0;1]$)
 7:   while $(k\leq n)\wedge(p>e^{-\lambda})$ do
-8:              $k$leftarrow $k+1$
-[…]
+8:     $k$ $\leftarrow$ $k+1$
+9:     $p$ $\leftarrow$ $p\cdot\text{{Unif}($[0;1]$)}$
 10:  until $k\leq n$
@@
 Note that this algorithm relies on real numbers arithmetic.
-In practice, approximating these numbers by IEEE 754 floating points numbers [ieee754] should introduce an acceptably small error.
+In practice, approximating these numbers by IEEE 754 floating points numbers [40] should introduce an acceptably small error.
 Indeed, since we only compute products (no sums or subtractions), which generally have few terms, the probability that they introduce an error should not be too far from $2^{-53}$ on a 64-bits architecture.
@@
 One possible way would be to use fixed points arithmetic for storing $p$ and to lazily generate the base 2 expansions of the uniform variables at play until we have enough bits to decide how $p$ and $e^{-\lambda}$ compare at line 7.
-Another way would be to use Ball arithmetic [Hoeven2010, arb] and to increase precision every time the comparison requires more bits.
+Another way would be to use Ball arithmetic [24, 26] and to increase precision every time the comparison requires more bits.
 The proofs of correctness and complexity below obviously assume such an implementation.
@@
 Proof.
-The correctness of Algorithm 4 follows from the arguments given in [Knuth1997, page 137], which we do not recall here.
-Regarding Algorithm 3, the for loop at line 6 implements the Fisher-Yates [FY1948] algorithm, which performs a uniform permutation of the contents of the array independently of its contents.
+The correctness of Algorithm 4 follows from the arguments given in [28, page 137], which we do not recall here.
+Regarding Algorithm 3, the for loop at line 6 implements the Fisher-Yates [14] algorithm, which performs a uniform permutation of the contents of the array independently of its contents.
 In our use-case, this implies that:
@@
 This tells us that, the probability that Algorithm 3 yields a particular variation with $p$ zeros is
-$\mathbb{P}[\textsc{BoundedPoisson}(1,n)=p]\cdot\frac{p!}{n!}=\frac{1}{p!\sum_{k=0}^{n}\frac{1}{k!}}\cdot\frac{p!}{n!}=\frac{1}{v_{n}}.\qed$ ∎
+$\mathbb{P}[\textsc{BoundedPoisson}(1,n)=p]\cdot\frac{p!}{n!}=\frac{1}{p!\sum_{k=0}^{n}\frac{1}{k!}}\cdot\frac{p!}{n!}=\frac{1}{v_{n}}.\qed$
 The minimal “amount of randomness” that is necessary to simulate a probability distribution is given by its entropy.
@@
 This can be done lazily, and again, it is sufficient to know a constant number of bits of these variables in average to decide whether $p>e^{-1}$.
-Regarding the shuffling happening at line 6 in Algorithm 3, it needs to draw $(n-1)$ uniform integers, respectively smaller or equal to 111122223333, …, $n-1$.
+Regarding the shuffling happening at line 6 in Algorithm 3, it needs to draw (n1)(n-1) uniform integers, respectively smaller or equal to 112233, …, n1n-1.
 At the first order, this incurs a total cost in terms of random bits, of
@@
 function UnifDOAGNaive($n$)
-     $A={(a_{i,j})}_{1\leq i,j\leq n}$leftarrow a zero-filled $n\times n$ matrix
+  $A={(a_{i,j})}_{1\leq i,j\leq n}$ $\leftarrow$ a zero-filled $n\times n$ matrix
   repeat
-         for $i$ from $1$ to $n-1$ do
-              ${(a_{i,j})}_{i<j\leq n}$leftarrow UnifVariation($n-i$)
-     until $A$ encodes a DOAG
-     return The DOAG corresponding to $A$
+   for $i$ from $1$ to n1n-1 do
+     ${(a_{i,j})}_{i<j\leq n}$ $\leftarrow$ UnifVariation(nin-i)
+  until AA encodes a DOAG
+  return The DOAG corresponding to AA
 Checking the validity of a matrix at line 8 corresponds to checking the conditions given in Theorem 3 at page 3.
@@
 This heuristic argument, hints at a more sparing algorithm that would start by filling the matrix near the diagonal and check its validity early, before generating the content of the whole matrix.
-This idea, of performing rejection as soon as possible in the generation process, is usually referred to as “anticipated rejection” and also appears in [DFLS2004] and [BPS1994] for instance.
+This idea, of performing rejection as soon as possible in the generation process, is usually referred to as “anticipated rejection” and also appears in [10] and [1] for instance.
 To put this idea in practice, we need to implement lazy variation generation, to be able to make progress in the generation of each line independently, and to perform the checks of Theorem 3 with as little information as necessary.
@@
 So a first ingredient of our optimised sampler is the following setup for lazy generation:
-  • for each row of the matrix (i.e. each variation to be sampled), we draw a Poisson variable $p_{i}$ of parameter $1$ and bounded by $(n-i)$;
+  • for each row of the matrix (i.e. each variation to be sampled), we draw a Poisson variable $p_{i}$ of parameter $1$ and bounded by (ni)(n-i);
   • drawing the number at position $(i,j)$, once we have drawn all the numbers of lower coordinate in the same row, can be done by selecting uniformly at random a cell with higher or equal coordinate on the same row and swapping their contents.
@@
 Instead, we interpret the numbers strictly larger than $(n-i-p_{i})$ as zeros.
-In this example $(n-i)=10$ and $p_{i}=3$ so the numbers 88889999, and $10$ are seen as zeros.
+In this example (ni)=10(n-i)=10 and $p_{i}=3$ so the numbers 8899, and $10$ are seen as zeros.
 Ingredient three: column by column checking
 The last detail that we need to explain is how to check the conditions of Theorem 3. As a reminder
-  • for each $2\leq j\leq n$ we need to compute the number $b_{j}=\max\,\{i\ |\ a_{i,j}>0\}$ (or 00 if this set is empty);
+  • for each $2\leq j\leq n$ we need to compute the number $b_{j}=\max\,\{i\ |\ a_{i,j}>0\}$ (or 0 if this set is empty);
   • we must check whether this sequence is weakly increasing;
@@
 For each column, we stop as soon as either a non-zero number is found or the constraints from Theorem 3 are violated.
-In order to check these constraints, while filling column $j$ from bottom $(i=j-1)$ to top, we halt as soon as either the cell on the left of the current cell, or the current cell is non-zero.
+In order to check these constraints, while filling column $j$ from bottom (i=j1)(i=j-1) to top, we halt as soon as either the cell on the left of the current cell, or the current cell is non-zero.
 The case when the left cell is non-zero corresponds to when $i=b_{j-1}$ and the conditions of Theorem 3 can be checked.
@@
 The second part of the algorithm, starting from line 23, completes the generation of the matrix once its near-diagonal part is known and we know no rejection is possible any more.
-This includes replacing some values of the matrix by 00 because of ingredient two above.
+This includes replacing some values of the matrix by 0 because of ingredient two above.
 Algorithm 6 An optimised uniform random sampler of DOAGs based on anticipated rejection
@@
 3:function UnifDOAGFast($n$)
-4:     $A={(a_{i,j})}_{1\leq i,j\leq n}$leftarrow the strictly upper triangular matrix ${(\boldsymbol{1}_{\{j>i\}}\cdot(j-i))}_{1\leq i,j\leq n}$
-5:     ${(p_{i})}_{1\leq i<n}$leftarrow uninitialised array
-6:     ${(s_{i})}_{1\leq i<n}$leftarrow uninitialised array
-7:     $\triangleright$ Anticipated rejection phase
-8:         $(i,j)$leftarrow $(1,2)$$\triangleright$ position of the current cell
-9:         $p_{1}$leftarrow BoundedPoisson($n-1$)
+4:  $A={(a_{i,j})}_{1\leq i,j\leq n}$ $\leftarrow$ the strictly upper triangular matrix ${(\boldsymbol{1}_{\{j>i\}}\cdot(j-i))}_{1\leq i,j\leq n}$
+5:  ${(p_{i})}_{1\leq i<n}$ $\leftarrow$ uninitialised array
+6:  ${(s_{i})}_{1\leq i<n}$ $\leftarrow$ uninitialised array
+[…]
+8:   $(i,j)$ $\leftarrow$ $(1,2)\triangleright$ position of the current cell
+9:   $p_{1}$ $\leftarrow$ BoundedPoisson(n1n-1)
 10:   while $j\leq n$ do
-11:              $r$leftarrow Unif($\llbracket j;n\rrbracket$)
+11:     $r$ $\leftarrow$ Unif($\llbracket j;n\rrbracket$)
 12:     $a_{i,r}\leftrightarrow a_{i,j}$
-13:              $s_{i}$leftarrow $j$
+13:     $s_{i}$ $\leftarrow$ $j$
 14:     if $(a_{i,j-1}\leq n-i-p_{i})\wedge(a_{i,j}\not\in\llbracket a_{i,j-1}+1;n-i-p_{i}\rrbracket)$ then
-15:                  $\triangleright$ Rejection
+[…]
 16:     else if $a_{i,j}\leq n-i-p_{i}$ then
-17:                  $j$leftarrow $j+1$
-18:                  $i$leftarrow $j-1$
-19:                  $p_{i}$leftarrow BoundedPoisson($1,n-i$)
+17:      $j$ $\leftarrow$ $j+1$
+18:      $i$ $\leftarrow$ j1j-1
+19:      $p_{i}$ $\leftarrow$ BoundedPoisson(1,ni1,n-i)
 20:     else
-21:                  $i$leftarrow $i-1$
+21:      $i$ $\leftarrow$ i1i-1
 22:  until $j>n$
-23:     for $i=1$ to $n-2$ $\triangleright$ Completion of the matrix
+[…]
 24:   for $j=i+1$ to $s_{i}$ do
@@
 26:   for $j=s_{i}+1$ to $n$ do
-27:              $r$leftarrow Unif($\llbracket j;n\rrbracket$)
+27:     $r$ $\leftarrow$ Unif($\llbracket j;n\rrbracket$)
 28:     $a_{i,r}\leftrightarrow a_{i,j}$
 29:     if $a_{i,j}>n-i-p_{i}$ then $a_{i,j}\leftarrow 0$
-30:     return $A$
+30:  return AA
 Lemma 8 (Correction of Algorithm 6).
@@
 In the rejection phase, in each column, we draw a certain number of zeros and at most one non-zero value before deciding whether to reject the matrix or to proceed to the next column.
-As a consequence, when lazily generating a variation matrix, we see at most $(n-1)$ non-zero values and a certain number of zeros that we can trivially upper-bound by the total number of zeros (strictly above the diagonal) in the matrix.
+As a consequence, when lazily generating a variation matrix, we see at most (n1)(n-1) non-zero values and a certain number of zeros that we can trivially upper-bound by the total number of zeros (strictly above the diagonal) in the matrix.
 The number of variations of size $n$ with exactly $p$ zeros (with $0\leq p\leq n$) is given by $\frac{n!}{p!}$ by Lemma 3. As a consequence, the expected number of zeros of a variation is given by
@@
 A common and powerful tool in combinatorics is the use of generating function to tackle not only asymptotic estimation, but also the convergence in distribution of parameters, such as the number of edges.
-The book [FS2009] is a reference on this topic.
-Classical approaches using ordinary, exponential, or even graphical ([PD2019]) generating functions fail in our context due to the super factorial behaviour of the number of DOAGs.
+The book [15] is a reference on this topic.
+Classical approaches using ordinary, exponential, or even graphical ([9]) generating functions fail in our context due to the super factorial behaviour of the number of DOAGs.
 It remains an open question whether it is possible to design a generating function approach to our objects, which could help obtaining finer estimates, not only over $D_{n}$, but also over the low of the number of edges.
@@
 We argue that the model of multi-edge DOAGs is natural, maybe even more so than that of DOAGs, since they encode (partially) compacted plane trees.
-Quantitative aspects of tree compaction, in particular the typical compression rate, has been studied in the past [FSS1990] in a general setting.
-However, the dual point of view that consists in studying already compacted structures directly is a more recent topic, see [EFW2021] and [GW2024] for instance.
+Quantitative aspects of tree compaction, in particular the typical compression rate, has been studied in the past [16] in a general setting.
+However, the dual point of view that consists in studying already compacted structures directly is a more recent topic, see [12] and [21] for instance.
 The class of multi-edge DOAGs generalises the classes studied in those two papers.
@@
 However we have already proved that, in the uniform model, they are connected with high probability since they have only one source with high probability.
-Moreover, since $D_{n}$ grows extremely fast, we can also foresee that a uniform DOAG of size $n$ with two connected components will typically have one tiny component of size $1$ and a big component of size $n-1$, and that the asymptotic estimations of such graphs is straightforward.
+Moreover, since $D_{n}$ grows extremely fast, we can also foresee that a uniform DOAG of size $n$ with two connected components will typically have one tiny component of size $1$ and a big component of size n1n-1, and that the asymptotic estimations of such graphs is straightforward.
 This implies that sampling a uniform connected DOAG with $n$ vertices is already possible, and efficient, by rejection and the question of their direct enumeration is thus mostly of mathematical interest.
@@
 Finally, it is also natural to wonder whether our successful approach at efficiently sampling DOAG applies to labelled DAGs.
-Of course, the asymptotics of DAGs is known [Robinson1973].
+Of course, the asymptotics of DAGs is known [36].
 But if a matrix encoding similar to ours is feasible, that is an encoding whose combinatorial properties are understood well enough to avoid introduction any bias, then it might be possible to devise an efficient, pre-computation-free, uniform sampler for DAGs.
@@
 Our recent ongoing work on this topic suggests that those paths can indeed be sampled efficiently, which will be investigated further in the near future.
-Notable is that the approach presented in [KM2015] can also be seen as a way to work with matrices while maintaining uniformity by using a combinatorial encoding using ordered integer partitions.
+Notable is that the approach presented in [31] can also be seen as a way to work with matrices while maintaining uniformity by using a combinatorial encoding using ordered integer partitions.
 A caveat however is that their approach still requires a costly pre-processing, which we seek to avoid using a rejection-based approach.
-[…]
+Acknowledgements
+We would like to thank one of the anonymous referees for pointing us at related work that we had missed in the field of Statistics and Bayesian inference, and appeared to have a strong connection with our line of work.
+This research received financial support from the LIA IFUMI and the ANR PPS project ANR-19-CE48-0014.